### v3

# SMI: An Information-Theoretic Metric for Predicting Model Knowledge Solely from Pre-Training Signals

Abstract
The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent. Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets. In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task indicative of a model’s internal knowledge. We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance. In response to these challenges, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models. Subsequently, we develop a multi-template QA evaluation framework incorporating paraphrased question variants. Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring any additional training. The experimental results demonstrate that SMI outperforms co-occurrence-based baselines, achieving R² > 0.75 on models with over one billion parameters. Theoretical analysis further reveals the marginal benefits of scaling model size and optimizing data, indicating that the upper limit of specific QA task accuracy is approximately 80%. Our project is available at https://github.com/yuhui1038/SMI.
1 Introduction
The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" (OpenAI, 2023). However, the report provides little detail on the methodology behind this prediction approach.
This technique is critical for the pre-training of large language models (LLMs). On the one hand, the cost of a single pre-training run can reach hundreds of millions of dollars, making re-pre-training a
Figure 1: Illustration of the three primary challenges addressed in this work, with the third being the most crucial.
highly challenging endeavor (Jiang et al., 2024; Yang et al., 2024; Dubey et al., 2024). The prediction technique, which allows for accurate performance forecasting on certain tasks based only on the pre-training signals, helps mitigate the resource waste. On the other hand, it can guide the construction of pre-training data that are more aligned with target tasks, optimizing resource usage (Gururangan et al., 2020; Longpre et al., 2024).
In this paper, we aim to predict performance in closed-book QA, a vital downstream task indicative of a model’s internal knowledge (Roberts et al., 2020). This setting is particularly well aligned with the nature of pre-training data and best reflects a model’s ability to answer fact-based questions. Meanwhile, we encounter three major challenges (Figure 1):
Challenge1: Limited access to and understanding of pre-training corpora. Currently, few open-source foundational LLMs fully disclose their pre-training datasets. Initiating pre-training from scratch is a costly endeavor, requiring the collection of extensive data and substantial computational resources. For instance, Meta’s technical report states that pre-training the 405B-parameter Llama 3 model used 15.6 trillion tokens and 16,000 H100 GPUs, with estimated costs exceeding 100 million dollars (Dubey et al., 2024).
Challenge2: Limitations of current evaluation methods for pre-trained models. Given the characteristics of closed-book QA tasks, we can evaluate a model’s accuracy on these tasks to assess the factual knowledge it has internalized during pretraining (Roberts et al., 2020; Wang et al., 2021). However, most evaluation methods face challenges, such as being overly sensitive to specific in-context examples and having coarse granularity in test data segmentation (Min et al., 2022; Srivastava et al., 2024). These issues make it difficult to accurately assess whether the pre-trained model has effectively retained specific knowledge.
Challenge3: Limitations of frequency-based metrics in predicting model performance. The model’s success on the target task relies on its ability to acquire world knowledge during pre-training, a process heavily shaped by the data (Allen-Zhu and Li, 2024a; Wang et al., 2024). Previous research suggests that models tend to acquire knowledge more effectively when it is frequently encountered within the pre-training data (Kandpal et al., 2023; Mallen et al., 2023; Allen-Zhu and Li, 2024b). However, these frequency-based methods does not take into account the specificity of knowledge, and thus only demonstrates general trends. For example, despite similar frequencies of occurrence, the model may struggle to pinpoint the location of Apple’s headquarters due to the pervasive nature of the subject. In contrast, identifying NVIDIA’s headquarters might be relatively straightforward. Furthermore, frequency-based methods do not consider the impact of model size.
Our work primarily focuses on effectively addressing these three challenges to predict model knowledge from pre-training signals, thereby estimating the model’s performance on downstream tasks without any additional training. Specifically:
First, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained LLMs. With comprehensive access to the full pre-training data, we are able to systematically trace and examine the knowledge that each model was exposed to during pre-training, enabling a detailed understanding of the relationship between pre-training data and downstream performance.
Second, inspired by the evaluation methodology for fine-tuned LLMs (Ye et al., 2024), we develop a robust multimodal evaluation approach. For each category of QA tasks, we design 20 semantically equivalent but syntactically diverse evaluation templates, enabling a more accurate and reliable assessment of the model’s knowledge retention.
Finally, we introduce SMI, an information-theoretic metric that linearly associates pre-training data characteristics, model size, and QA accuracy without requiring additional training. Experimental results on 24 pre-trained models indicate that the SMI metric effectively predicts a model’s knowledge retention. Across various model sizes, ranging from 14 million to 176 billion parameters, SMI significantly outperforms the co-occurrence-based baseline. For models with more than 1 billion parameters, the coefficient of determination (R²) between SMI and QA accuracy exceeds 0.75, indicating a strong linear relationship.
Moreover, the comparative analysis of experimental results from different models further reveals that both scaling model size and optimizing data present diminishing returns. As the model size increases, its impact on performance in QA tasks becomes progressively smaller. Meanwhile, even when the SMI metric reaches its maximum on the largest models within each families, the accuracy of the QA tasks cannot surpass a ceiling of approximately 80%.
Overall, our contributions are threefold:
  1. 1.
    We propose an information theory-based method and introduce the SMI metric, which accurately predicts the knowledge retention of pre-trained LLMs using only pre-training signals, without requiring any additional training.
  2. 2.
    Across LLMs with varying sizes from 14 million to 176 billion parameters, the SMI metric exhibits a strong linear relationship with QA accuracy, significantly outperforms the traditional co-occurrence-based baseline.
  3. 3.
    Experimental results across diverse models reveal that both scaling model size and optimizing data yield diminishing returns. Even under optimal conditions, the upper bound of accuracy on the evaluated QA tasks remains around 80%.
2 Related Work
Pre-training data and knowledge retention
The characteristics of pre-training data are crucial for the knowledge retention and downstream task performance of LLMs. Numerous studies highlight the significance of repetition frequency in this process. For instance, Carlini et al. (2023) identified a logarithmic relationship between repetition frequency and memory effects, while Chowdhery et al. (2023) demonstrated that sequences repeated over 500 times are completed by models with an accuracy exceeding 40%. Similarly, Ju et al. (2024) investigated the influence of data frequency on multi-hop reasoning, uncovering the presence of “factual shortcuts”. Allen-Zhu and Li (2024b) proposed that exposing knowledge 1000 times can achieve a storage capacity of two bits per parameter in LLMs. Additionally, Razeghi et al. (2022) and Yadlowsky et al. (2023) demonstrated that lower-order co-occurrences in pre-training data enhance numerical reasoning, while McCoy et al. (2023) associated the prevalence of tasks in pre-training data with better performance on tasks like ROT13.
Model size and knowledge retention
Furthermore, the model size significantly influences the knowledge retention and capabilities of LLMs. Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern (Kaplan et al., 2020).  Kandpal et al. (2023) evaluate the BLOOM model families (Scao et al., 2022) on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters. Additionally, the memory capacity of a model closely correlates with its size (Li et al., 2024).
Knowledge triples as an evaluation tool
Knowledge triples (subject, relation, object) serve as an effective tool for evaluating the knowledge retention of LLMs. Petroni et al. (2019) employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks. Subsequently, He et al. (2024); Ju et al. (2024) further explore the application of knowledge triples in reasoning tasks. Moreover, Wang et al. (2024) and Allen-Zhu and Li (2024a) use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
3 Methodology
To address the three challenges outlined in Section 1, we propose a method to predict model knowledge solely from pre-training signals. We begin with prior work and explore the use of the co-occurrence metric. However, the co-occurrence metric do not account for the specificity of knowledge. Inspired by information theory, we propose the Mutual Information (MI) metric that simultaneously considers both the frequency and specificity of knowledge, significantly enhancing prediction outcomes. Nevertheless, the MI metric does not consider the impact of model size, which has been shown to correlate closely with model knowledge retention. Finally, we propose the SMI metric, a composite representation of pre-training data characteristics and model size.
3.1 Starting from the Co-occurrence Metric
Co-occurrence metric baseline
Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns (Kandpal et al., 2023; Mallen et al., 2023; Allen-Zhu and Li, 2024b). Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach (Kandpal et al., 2023). There is a linear relationship between QA accuracy and the logarithm of the number of documents (Kandpal et al., 2023). The logarithm of the document count serves as the co-occurrence metric, which constitutes our baseline.
Knowledge triples
In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object (Wang et al., 2024). For a given LLM, denoted as $F$, we define $F$ as mastering the knowledge triple $t$, thus demonstrating competence in the corresponding QA task, as follows:
$F(q_{s,r})=o,\quad(q\in\mathbf{Q},o\in\mathbf{O}).$ (1)
Here, $q_{s,r}$ represents the combined representation of the subject $s$ and the relation $r$, while $o$ denotes the representation of the object. For instance, if $s=\text{Apple}$ and $r=\text{headquarter}$, then $q_{s,r}$ can be expressed as statements like “The headquarters of Apple is in…”, “Apple is headquartered in…”, or “Apple’s head office is based in…”. Similarly, $o$ corresponds to object representation, such as “Cupertino”, “Cupertino, California”, or “Cupertino city”. The sum of all $q_{s,r}$ representations is denoted as $\mathbf{Q}$, and the sum of all $o$ representations is denoted as $\mathbf{O}$.
Multi-template evaluation method
Evaluating the knowledge retention of LLMs presents challenges due to the diverse range of queries formed by the subject and relation of knowledge triples. Inspired by the evaluation method for fine-tuned LLMs (Ye et al., 2024), we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
Specifically, by constructing synonymous sentences, we generate a substantial number of query templates that are semantically similar but vary in form. For each type of knowledge triple, we produce 20 templates that differ in length and structure. Each template $q$ represents a specific instance within the query set $\mathbf{Q}$, and these 20 selected $q$ templates approximate the entirety of $\mathbf{Q}$.
Linear relationship
On the one hand, we calculate metrics from the pre-training data. Each QA task consists of a subject and an object. We perform document-level retrieval on the pre-training data and compute three values: $P(s)$ represents the proportion of documents containing the subject out of the total number of documents $\mathbf{N}$ in the corpus, $P(o)$ denotes the proportion of documents containing the object out of $\mathbf{N}$, and $P(s,o)$ signifies the proportion of documents containing both the subject and the object out of $\mathbf{N}$. The co-occurrence metric is given by $\log_{2}(P(s,o))$.
On the other hand, using the multi-template evaluation method, we obtain the model’s accuracy on QA tasks. We fit a linear equation between the co-occurrence metric and QA accuracy, and then calculate R² and Mean Square Error (MSE) to assess the strength and stability of their linear relationship.
3.2 Considering the Knowledge Specificity to Obtain the MI Metric
Knowledge specificity
The co-occurrence metric does not account for the specificity of knowledge. When entities of knowledge are common phrases that frequently appear in pre-training data, their specificity is low, making the knowledge more difficult to retain. Conversely, if the entities appear infrequently in pre-training data, their specificity is high, and the knowledge is easier to retain. For example, the address of NVIDIA’s headquarters is easier to remember than that of Apple’s headquarters, because the word "Apple" occurs more frequently across diverse contexts. MI in information theory considers both the co-occurrence of two variables and their individual specificities (Shannon, 1948). Therefore, we aim to use MI to improve the co-occurrence metric.
MI metric
Unlike the MI in information theory, which considers all possible values of random variables, the MI metric within pre-training data considers only those instances in which the relevant entities actually appear. Given pre-training data $\mathbf{P}$ consisting of $\mathbf{N}$ paragraphs, a knowledge triple $(s,r,o)$, and corresponding $P(s)$, $P(o)$, and $P(s,o)$, the MI metric between $s$ and $o$ is defined as:
$I(s,o)$ $=P(s,o)\log\left(\frac{P(s,o)}{P(s)P(o)}\right),$ (2)
$MI(s,o)$ $=\log_{2}(I(s,o)).$ (3)
In the above formula, $P(s)$ and $P(o)$ penalize the number of documents containing the subject and object, respectively, to emphasize the specificity of $s$ and $o$. The term $I(s,o)$ represents the amount of shared information between $s$ and $o$, and it can also indicate the reduction in the amount of information about $o$ given $s$. During inference with LLMs, this amount of information refers to the probability of the model generating the object given the subject. Considering the uneven distribution of $I(s,o)$, we take its logarithm to obtain the MI metric (Figure 2).
3.3 Considering the Model Size to Obtain the Size-dependent MI metric
Model size
The MI metric does not account for the impact of model size. Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger (Kaplan et al., 2020). In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters (Kandpal et al., 2023). Therefore, we propose to enhance the MI metric by incorporating model size $\Phi$, measured in billions of parameters.
The method of integrating model size is not unique. A guiding principle is that, for the same data, a higher $\Phi$ should correspond to a higher value of the new metric, reflecting improved QA performance. However, the gain in accuracy should diminish progressively. Importantly, as $\Phi$ approaches infinity, the new metric should converge to the MI metric, indicating that the model’s memory capacity is sufficient to fully capture the relevant knowledge.
Size-dependent MI metric
Based on these principles, we propose the SMI metric—an exponential function with the normalized MI as the base and $1+\frac{1}{\Phi}$ as the exponent.
$SMI(s,o,\Phi)=Norm(MI(s,o))^{1+\frac{1}{\Phi}}.$ (4)
In the above formula, the MI metric quantifies the information of knowledge triples in the pre-training data, while the model size $\Phi$ implies the memory capacity of LLMs. These two factors jointly determine the model’s knowledge retention. We normalize the MI metric from its original range of -30 to 0 onto a 0–1 scale. Since the normalized MI metric is always less than 1, the factor $1+\frac{1}{\Phi}$ decreases as $\Phi$ increases, leading to a larger SMI value, which indicates stronger knowledge retention.
Figure 2: Evaluation data statistics. Left: Co-occurrence frequencies of QA entity pairs in The Pile corpus. Right: Distribution of 12,671 QA tasks across 15 relation types.
4 Experiments
First, we clean and enhance an open-source Wikipedia knowledge triple dataset to make it suitable for inference scenarios using pre-trained models. Next, we conduct retrieval analysis on five large-scale open-source pre-trained corpora. Finally, we perform experimental analysis on 21 open-source pre-trained models and 3 models that we pre-train ourselves.
4.1 Experimental Setup
Evaluation data
We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours (Elazar et al., 2021). We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2. For each category of knowledge triple relations, we create 20 synonymous query templates to construct the QA task evaluation dataset.
Pre-training corpora
We conduct a document-level retrieval analysis of the following pre-training datasets, selected based on their usage in the LLMs under our consideration:
  • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs. It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality (Gao et al., 2021).
  • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project. It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency (Laurençon et al., 2022; Scao et al., 2022).
  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs (Soboleva et al., 2023).
  • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes. It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM (Penedo et al., 2023).
  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency (Foundation, ).
Pre-training data retrieval
We process the pre-training corpus at the document level. For each knowledge triple, we retrieve the frequency of occurrences of the subject, the object, and their co-occurrences. Using the retrieval results from all documents, we calculate the co-occurrence metric and MI metric for each knowledge triple. Subsequently, we compute the SMI metric based on the MI metric and the model size. The retrieval process runs on 100 CPUs for 3 days.
Pre-trained models
We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs (Vaswani et al., 2017). These models not only make the model weights available but also provide complete pre-training data. They are primarily categorized into the following families:
  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion (Scao et al., 2022).
  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile (Black et al., 2021, 2022; Wang, 2021).
  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters. These models use a uniform data and training process, facilitating research that is both replicable and interpretable (Biderman et al., 2023).
  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata. It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions (Zhang et al., 2024; Soboleva et al., 2023; Li et al., 2023). The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
Additionally, we pre-train three models, named ours-1.6b, ours-7b, and ours-13b, based on RefinedWeb and Wikipedia as a supplement to the aforementioned models.
Evaluation setting
We assess these pre-trained models using the designated evaluation dataset. During the evaluation process, we set the temperature to 0.7, the maximum new tokens to 32, and the precision to bfloat16. We consider a response to be correct if the object appears within it. To ensure stability in the evaluation, we perform inference 20 times for each of the 20 queries in the QA task, subsequently calculating the average accuracy of 400 responses. This inference process runs for 5 days on 8 H20 GPUs.
The results of the pre-training data retrieval indicate that the co-occurrence frequency distribution of question entities and answer entities is highly uneven (Figure 2). The majority of QA data co-occur between 10 and 1000 times. In contrast, QA data with higher co-occurrence frequencies are notably sparse. To ensure data balance and effectively evaluate data across different ranges, we group QA data with similar metric values. For co-occurrence metrics and MI metrics, we calculate the average accuracy using intervals of 0.2. The SMI metric is computed based on the MI metric.
Finally, we perform linear regression on the metrics and accuracy, and calculate the R² to assess the strength of the linear relationship. Simultaneously, to evaluate the prediction error of the regression equation in specific QA tasks, we compute the MSE for each QA task and take the average.
Table 1: R² and MSE (×$\times 10^{-2}$) results of the CO-OCCUR, MI, and SMI metrics across different models, ranging from 14M to 176B parameters. Across all models, MI consistently outperforms CO-OCCUR, and SMI further improves upon MI.
Model R² ($\uparrow$ higher is better) MSE (×$\times 10^{-2}$, $\downarrow$ lower is better)
CO-OCCUR MI SMI CO-OCCUR MI SMI
BLOOM
bloom-560m 0.384 0.589 0.624 8.566 7.729 7.652
bloom-1b1 0.577 0.715 0.756 9.123 7.950 7.824
bloom-1b7 0.584 0.740 0.772 9.389 8.083 7.970
bloom-3b 0.705 0.782 0.807 9.314 7.863 7.764
bloom-7b1 0.676 0.804 0.814 9.943 8.309 8.263
bloom (176b) 0.644 0.764 0.764 7.327 6.029 6.029
GPT-Neo
gpt-neo-125m 0.349 0.531 0.593 7.004 7.107 6.999
gpt-neo-1.3B 0.671 0.810 0.859 8.233 7.444 7.265
gpt-neo-2.7B 0.672 0.826 0.850 8.168 7.320 7.233
gpt-j-6b 0.631 0.862 0.869 9.026 7.720 7.692
gpt-neox-20b 0.593 0.841 0.843 8.557 7.476 7.468
Pythia
pythia-14m 0.240 0.375 0.496 6.855 7.183 6.978
pythia-31m 0.287 0.381 0.549 6.905 7.202 6.905
pythia-70m 0.355 0.430 0.585 6.749 7.051 6.783
pythia-160m 0.454 0.522 0.611 6.877 6.969 6.796
pythia-410m 0.524 0.701 0.759 7.439 7.035 6.889
pythia-1.4b 0.574 0.799 0.835 8.254 7.325 7.202
pythia-2.8b 0.631 0.845 0.866 8.593 7.505 7.421
pythia-6.9b 0.637 0.855 0.865 8.538 7.373 7.334
pythia-12b 0.602 0.853 0.858 8.741 7.514 7.494
TinyLlama
TinyLlama-1.1B 0.641 0.760 0.839 7.803 7.135 6.887
Ours
ours-1.6b 0.710 0.794 0.848 8.563 7.696 7.462
ours-7b 0.784 0.847 0.855 8.983 8.033 7.997
ours-13b 0.661 0.816 0.820 8.585 7.688 7.674
(a) $\text{R}^{2}$ on smaller models
(b) $\text{R}^{2}$ on larger models
Figure 3: Comparison of $\text{R}^{2}$ results for three metrics—CO-OCCUR, MI, and SMI—across models of different sizes. (a) Smaller models (bloom-560m, gpt-neo-1.3B, pythia-14m, ours-1.6b); (b) Larger models (bloom, gpt-neox-20b, pythia-12b, ours-13b). $\uparrow$ Higher is better.
(a) bloom (176b)
(b) gpt-neox-20b
(c) pythia-12b
(d) ours-13b
Figure 4: Relationship between the SMI metric and average accuracy on QA tasks for different model series: (a) BLOOM (176B), (b) GPT-Neo (20B), (c) Pythia (12B), and (d) Ours (13B).
Table 2: Hyperparameter analysis of the SMI metric across different BLOOM models. (a) illustrates the impact of varying unit scales, while (b) shows the effect of different pre-normalization ranges.
(a) SMI values across varying unit scales
Unit bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) $10^{8}$ 0.590 0.716 0.741 0.783 0.804 0.764 $10^{9}$ 0.600 0.723 0.745 0.785 0.805 0.764 $10^{10}$ 0.624 0.756 0.772 0.807 0.814 0.764 $10^{11}$ 0.394 0.497 0.586 0.760 0.826 0.766 $10^{12}$ 0.343 0.222 0.197 0.235 0.349 0.769
(b) SMI values under different pre-normalization ranges
Range bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) $[-30,0]$ 0.624 0.756 0.772 0.807 0.814 0.764 $[-35,0]$ 0.626 0.750 0.766 0.800 0.811 0.764 $[-40,0]$ 0.625 0.746 0.761 0.797 0.809 0.764 $[-45,0]$ 0.623 0.742 0.758 0.794 0.808 0.764 $[-50,0]$ 0.620 0.739 0.756 0.793 0.808 0.764
4.2 Experimental Results
Results of SMI, MI, and co-occurrence metrics
The primary experimental results for the three metrics are presented in Table 1. Overall, we observe the following: (1) The SMI metric performs
Figure 5: Relationship between average QA accuracy on the full evaluation dataset and model size across five different model families. Model size is reported in billions of parameters and plotted on a base-2 logarithmic scale. Variations in pre-training data across model families may account for differences in performance.
exceptionally well on the evaluation set, with both R² and MSE surpassing the baseline across all models. (2) For all models, the SMI metric achieves R² values greater than 0.49, and for models with parameters exceeding 1 billion, the R² exceeds 0.75. This highlights a robust linear relationship and effectively predicts QA accuracy. (3) The MI metric outperforms the baseline co-occurrence metric, while the SMI metric further exceeds the MI metric, demonstrating the effectiveness of our iterative improvement approach.
The comparison of experimental results from models of different sizes indicates diminishing returns when increasing model size. A linear relationship is observed between the models’ average accuracy on the full evaluation dataset and the logarithm of model size, suggesting that significant gains in QA performance require exponential increases in model size (Figure 5). Meanwhile, the advantage of the SMI metric over the MI metric is more pronounced in smaller models (Figure 3). As model size increases, this advantage becomes less evident. For the BLOOM 560m model, the R² values differ by 3.55E-02, and the MSE values differ by 7.75E-04. In contrast, for the BLOOM 176b model, the difference in R² values is only 2.42E-04, and the difference in MSE values is just 6.60E-06. This indicates that larger models may possess sufficient memory capacity, thereby diminishing the impact of further scaling.
Moreover, the benefits of data optimization exhibit diminishing returns and an upper limit. We find that although the models’ performance on the QA task improves with increasing SMI, the accuracy remains relatively low even when the SMI reaches its maximum value (see Figure 4). Specifically, the upper accuracy limits for the models bloom, gpt-neox-20b, pythia-12b, and ours-13b are 0.638, 0.727, 0.75, and 0.771, respectively. This indicates that even with perfect optimization of the pre-training data for these models, the accuracy on the QA task cannot exceed an approximate upper limit of 0.8.
Results of hyperparameter influence on the SMI metric
The SMI framework involves two critical hyperparameters. The first is the unit for model parameter quantity, which is set by default to 1e9. The second is the pre-normalization range for the SMI metric, with a default setting of -30 to 0. We conduct ablation experiments on these two hyperparameters, and the results are presented in Table 2. From the experimental results, we observe the following: (1) The unit for model parameter quantity significantly affects the SMI results. Using 1e10 as the unit yields the best overall performance. However, as the model size increases, the optimal unit value also increases. Smaller models exhibit lower robustness to parameter variations, whereas larger models display enhanced robustness. (2) The pre-normalization range has a relatively minor effect on the SMI results. The range of -30 to 0 performs best overall. Although a smaller pre-normalization range is preferable, the minimum value in the experimental data approximates -30, suggesting that further reduction might result in data loss.
5 Conclusion
In this paper, we explore how to predict model knowledge using only pre-training signals and propose effective solutions to address three major challenges. We propose SMI, an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy without any additional training. SMI achieves R² > 0.75 on models with over one billion parameters. Experimental results further reveal that both scaling model size and optimizing data provide marginal benefits, with the accuracy on specific QA tasks plateauing at approximately 80%.
This paper emphasizes the feasibility of predicting downstream performance based on pre-training signals. It offers a principled metric for data-model-task alignment, thereby providing a reference for optimizing data selection and performance prediction during the pre-training phase of LLMs. To facilitate further research, all evaluation data and analysis code are made publicly available.
6 Limitations
Although the SMI metric performs well in predicting the accuracy of pretrained models on specific QA tasks, the prediction results remain highly dependent on the type of task. The relationship between the SMI and performance may vary across different categories of tasks, necessitating further investigation into various types of tasks in the future.
Additionally, the fitting results for models with differing parameter sizes exhibit slight differences, and smaller models demonstrate increased sensitivity to the hyperparameters of the SMI. Future research will focus on developing representations of pretrained signals that are more robust to model size and exploring the mathematical foundations underlying these representations.
Appendix A Technical Appendices and Supplementary Material
A.1 Licenses and Terms of Use
We use the following publicly available assets in this work:
We confirm that all licenses and terms of use are respected.
A.2 Templates for Knowledge Triples
Table 3: Twenty templates and their corresponding accuracy values for the $\mathbf{BORN}$ relation on the ours-13b model. [S] represents the subject in the knowledge triple.
Template ACC
[S]’s birth took place in 0.186
[S] was brought into existence in 0.187
[S] was born into the world in 0.190
[S]’s nativity is 0.197
The place in which [S] was given birth to is 0.201
The hometown of [S] is 0.203
The birthplace of [S] is 0.212
[S] was delivered in 0.215
[S]’s cradle was in 0.219
[S]’s entry into the world happened in 0.225
[S] first saw the light in 0.228
[S]’s roots are in 0.230
[S]’s life began in 0.230
[S] was born and raised in 0.237
[S]’s origin is 0.238
[S] entered life in 0.251
[S] came into the world in 0.255
[S] was born in 0.255
[S]’s birth occurred in 0.263
[S]’s starting point was 0.282
[…]
A.3 Case Study
Table 4: Several cases with significant deviations from the co-occurrence metric are shown here. $N_{s}$ is the number of paragraphs in which the subject appears, $N_{o}$ is the number of paragraphs in which the object appears, and $N_{s,o}$ is the number of paragraphs in which they co-occur. The first two cases exhibit rare co-occurrence but demonstrate high accuracy, whereas the latter two cases show frequent co-occurrence but exhibit low accuracy.
Subject Question Object $N_{s}$ $N_{o}$ $N_{s,o}$ $ACC$
Rare Co-occurrence but High Accuracy
Csepel SC The headquarters of Csepel SC is in Budapest 398 748,860 172 0.655
Liis Lemsalu Liis Lemsalu is a citizen of Estonia 307 967,207 160 0.570
Frequent Co-occurrence but Low Accuracy
Paris Paris is located in Europe 20,902,571 59,408,155 3,761,610 0.043
Ohio The capital of Ohio is Columbus 11,894,303 3,718,483 1,210,957 0.193
Table 4 highlights four counterintuitive cases in our pre-training data. From these, we observe the following: (1) In the first two cases, the subject and object rarely co-occur, suggesting a seemingly weak relationship. However, whenever the subject does appear, the object almost always appears alongside it, resulting in high accuracy. (2) In the last two cases, despite frequent co-occurrence between the subject and object, their individual occurrence frequencies are significantly higher, leading to low accuracy. For example, when prompted with “Paris is located in”, the model is more likely to respond with “France” rather than “Europe”.
These examples highlight the pivotal role of knowledge specificity in the memory of LLMs. It is not simply the frequencies of knowledge occurrence that enhance retention, but rather the specificity of the knowledge that proves more influential. This insight suggests that in downstream task training, even when domain-specific data is significantly fewer than the pre-training data, minimizing repetition with the pre-training data can increase the SMI, making it easier for LLMs to retain the relevant knowledge.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
[…]
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
[…]
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
[…]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
[…]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
[…]
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
[…]
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
[…]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
[…]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
[…]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A] This paper constitutes fundamental research and is not associated with any specific applications or deployments.
[…]
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
[…]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
[…]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
[…]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
[…]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
[…]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
[…]
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v4

@@
-SMI: An Information-Theoretic Metric for Predicting Model Knowledge Solely from Pre-Training Signals
+Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training
 Abstract
@@
 Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets.
-In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task indicative of a model’s internal knowledge.
+In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools.
 We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance.
-In response to these challenges, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models.
-Subsequently, we develop a multi-template QA evaluation framework incorporating paraphrased question variants.
-Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring any additional training.
-The experimental results demonstrate that SMI outperforms co-occurrence-based baselines, achieving R² > 0.75 on models with over one billion parameters.
-Theoretical analysis further reveals the marginal benefits of scaling model size and optimizing data, indicating that the upper limit of specific QA task accuracy is approximately 80%.
-Our project is available at https://github.com/yuhui1038/SMI.
+In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models.
+We then develop a multi-template QA evaluation framework incorporating paraphrased question variants.
+Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training.
+Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters.
+Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
+Abstract
+The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent.
+Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets.
+In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools.
+We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance.
+In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models.
+We then develop a multi-template QA evaluation framework incorporating paraphrased question variants.
+Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training.
+Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters.
+Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
+[…]
+[…]
+Figure 1: Illustration of the three primary challenges addressed in this work, with the third being the most crucial.
 1 Introduction
-The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" (OpenAI, 2023).
+The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" [25].
 However, the report provides little detail on the methodology behind this prediction approach.
 This technique is critical for the pre-training of large language models (LLMs).
-On the one hand, the cost of a single pre-training run can reach hundreds of millions of dollars, making re-pre-training a
-Figure 1: Illustration of the three primary challenges addressed in this work, with the third being the most crucial.
-highly challenging endeavor (Jiang et al., 2024; Yang et al., 2024; Dubey et al., 2024).
+On the one hand, the cost of a single pre-training run can reach hundreds of millions of dollars, making re-pre-training a highly challenging endeavor [14, 39, 8].
 The prediction technique, which allows for accurate performance forecasting on certain tasks based only on the pre-training signals, helps mitigate the resource waste.
-On the other hand, it can guide the construction of pre-training data that are more aligned with target tasks, optimizing resource usage (Gururangan et al., 2020; Longpre et al., 2024).
-In this paper, we aim to predict performance in closed-book QA, a vital downstream task indicative of a model’s internal knowledge (Roberts et al., 2020).
+On the other hand, it can guide the construction of pre-training data that are more aligned with target tasks, optimizing resource usage [12, 21].
+In this paper, we aim to predict performance in closed-book QA, a vital downstream task indicative of a model’s internal knowledge [29].
 This setting is particularly well aligned with the nature of pre-training data and best reflects a model’s ability to answer fact-based questions.
@@
 Challenge1: Limited access to and understanding of pre-training corpora.
-Currently, few open-source foundational LLMs fully disclose their pre-training datasets.
-Initiating pre-training from scratch is a costly endeavor, requiring the collection of extensive data and substantial computational resources.
-For instance, Meta’s technical report states that pre-training the 405B-parameter Llama 3 model used 15.6 trillion tokens and 16,000 H100 GPUs, with estimated costs exceeding 100 million dollars (Dubey et al., 2024).
+Currently, few open-source foundation models fully disclose their pre-training datasets.
+Pre-training from scratch is prohibitively expensive, requiring massive data collection and compute.
+For example, Meta reports that pre-training the 405B-parameter Llama 3 model used 15.6 trillion tokens and 16,000 H100 GPUs, with estimated costs over 100 million dollars [8].
 Challenge2: Limitations of current evaluation methods for pre-trained models.
-Given the characteristics of closed-book QA tasks, we can evaluate a model’s accuracy on these tasks to assess the factual knowledge it has internalized during pretraining (Roberts et al., 2020; Wang et al., 2021).
-However, most evaluation methods face challenges, such as being overly sensitive to specific in-context examples and having coarse granularity in test data segmentation (Min et al., 2022; Srivastava et al., 2024).
+Given the characteristics of closed-book QA tasks, we can evaluate a model’s accuracy on these tasks to assess the factual knowledge it has internalized during pretraining [29, 36].
+However, most evaluation methods face challenges, such as being overly sensitive to specific in-context examples and having coarse granularity in test data segmentation [24, 33].
 These issues make it difficult to accurately assess whether the pre-trained model has effectively retained specific knowledge.
+In particular, assessing a model’s knowledge using a single query template is unreliable [40].
 Challenge3: Limitations of frequency-based metrics in predicting model performance.
-The model’s success on the target task relies on its ability to acquire world knowledge during pre-training, a process heavily shaped by the data (Allen-Zhu and Li, 2024a; Wang et al., 2024).
-Previous research suggests that models tend to acquire knowledge more effectively when it is frequently encountered within the pre-training data (Kandpal et al., 2023; Mallen et al., 2023; Allen-Zhu and Li, 2024b).
+The model’s success on the target task relies on its ability to acquire world knowledge during pre-training, a process heavily shaped by the data [1, 37].
+Previous research suggests that models tend to acquire knowledge more effectively when it is frequently encountered within the pre-training data [16, 22, 2].
 However, these frequency-based methods does not take into account the specificity of knowledge, and thus only demonstrates general trends.
@@
 With comprehensive access to the full pre-training data, we are able to systematically trace and examine the knowledge that each model was exposed to during pre-training, enabling a detailed understanding of the relationship between pre-training data and downstream performance.
-Second, inspired by the evaluation methodology for fine-tuned LLMs (Ye et al., 2024), we develop a robust multimodal evaluation approach.
-For each category of QA tasks, we design 20 semantically equivalent but syntactically diverse evaluation templates, enabling a more accurate and reliable assessment of the model’s knowledge retention.
-Finally, we introduce SMI, an information-theoretic metric that linearly associates pre-training data characteristics, model size, and QA accuracy without requiring additional training.
+Second, inspired by the evaluation methods for fine-tuning LLMs [40], we develop a robust multimodal evaluation approach.
+Recognizing that the same knowledge can correspond to various questioning styles, each style allows for the design of corresponding templates.
+For each category of question-answering tasks, we design 20 semantically equivalent but syntactically diverse evaluation templates, enabling a more accurate and reliable assessment of the model’s knowledge retention.
+Finally, we propose SMI, an information-theoretic metric that linearly associates pre-training data characteristics, model size, and QA accuracy without requiring additional training.
 Experimental results on 24 pre-trained models indicate that the SMI metric effectively predicts a model’s knowledge retention.
@@
 For models with more than 1 billion parameters, the coefficient of determination (R²) between SMI and QA accuracy exceeds 0.75, indicating a strong linear relationship.
-Moreover, the comparative analysis of experimental results from different models further reveals that both scaling model size and optimizing data present diminishing returns.
-As the model size increases, its impact on performance in QA tasks becomes progressively smaller.
-Meanwhile, even when the SMI metric reaches its maximum on the largest models within each families, the accuracy of the QA tasks cannot surpass a ceiling of approximately 80%.
-Overall, our contributions are threefold:
+Moreover, the comparative analysis across models shows diminishing returns from both scaling model size and optimizing data.
+As model size increases, its marginal impact on QA performance decreases, suggesting that larger capacity alone cannot guarantee continued gains.
+Likewise, while higher SMI values are generally associated with better accuracy, pushing SMI further becomes increasingly difficult due to its logarithmic growth.
+Notably, even in the largest models within each family, QA accuracy consistently plateaus around 80% when SMI is maximized.
+This suggests a fundamental limitation in how much factual knowledge can be internalized during pre-training, highlighting the need for external augmentation methods such as retrieval or prompting in downstream stages.
+In summary, our contributions are threefold:
   1. 1.
@@
     Experimental results across diverse models reveal that both scaling model size and optimizing data yield diminishing returns.
-Even under optimal conditions, the upper bound of accuracy on the evaluated QA tasks remains around 80%.
+Even under optimal conditions, QA accuracy plateaus around 80%, reflecting the intrinsic limit of knowledge that can be memorized during pre-training.
 2 Related Work
 Pre-training data and knowledge retention
-The characteristics of pre-training data are crucial for the knowledge retention and downstream task performance of LLMs.
-Numerous studies highlight the significance of repetition frequency in this process.
-For instance, Carlini et al.
-(2023) identified a logarithmic relationship between repetition frequency and memory effects, while Chowdhery et al.
-(2023) demonstrated that sequences repeated over 500 times are completed by models with an accuracy exceeding 40%.
-Similarly, Ju et al.
-(2024) investigated the influence of data frequency on multi-hop reasoning, uncovering the presence of “factual shortcuts”.
-Allen-Zhu and Li (2024b) proposed that exposing knowledge 1000 times can achieve a storage capacity of two bits per parameter in LLMs.
+The characteristics of pre-training data are crucial for knowledge retention and downstream performance in LLMs.
+Many studies emphasize the role of repetition frequency.
+For example, Carlini et al.
+[6] identified a logarithmic relationship between repetition and memory effects, while Chowdhery et al.
+[7] showed that sequences repeated over 500 times are completed with over 40% accuracy.
+Ju et al.
+[15] found that high-frequency data can lead to “factual shortcuts” in multi-hop reasoning.
+Allen-Zhu and Li [2] suggested that exposing knowledge 1000 times enables a storage capacity of two bits per parameter.
 Additionally, Razeghi et al.
-(2022) and Yadlowsky et al.
-(2023) demonstrated that lower-order co-occurrences in pre-training data enhance numerical reasoning, while McCoy et al.
-(2023) associated the prevalence of tasks in pre-training data with better performance on tasks like ROT13.
+[28] and Yadlowsky et al.
+[38] linked lower-order co-occurrences to improved numerical reasoning, while McCoy et al.
+[23] associated task prevalence in pre-training with better performance on tasks like ROT13.
 Model size and knowledge retention
-Furthermore, the model size significantly influences the knowledge retention and capabilities of LLMs.
-Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern (Kaplan et al., 2020).
+Furthermore, the model size significantly influences the knowledge retention of LLMs.
+Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern [17].
 Kandpal et al.
-(2023) evaluate the BLOOM model families (Scao et al., 2022) on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters.
-Additionally, the memory capacity of a model closely correlates with its size (Li et al., 2024).
+[16] evaluate the BLOOM model families [30] on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters.
+Additionally, the memory capacity of a model closely correlates with its size [19].
 Knowledge triples as an evaluation tool
@@
 Petroni et al.
-(2019) employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks.
+[27] employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks.
 Subsequently, He et al.
-(2024); Ju et al.
-(2024) further explore the application of knowledge triples in reasoning tasks.
+[13], Ju et al.
+[15] further explore the application of knowledge triples in reasoning tasks.
 Moreover, Wang et al.
-(2024) and Allen-Zhu and Li (2024a) use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
+[37] and Allen-Zhu and Li [1] use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
 3 Methodology
@@
 We begin with prior work and explore the use of the co-occurrence metric.
-However, the co-occurrence metric do not account for the specificity of knowledge.
 Inspired by information theory, we propose the Mutual Information (MI) metric that simultaneously considers both the frequency and specificity of knowledge, significantly enhancing prediction outcomes.
-Nevertheless, the MI metric does not consider the impact of model size, which has been shown to correlate closely with model knowledge retention.
 Finally, we propose the SMI metric, a composite representation of pre-training data characteristics and model size.
@@
 Co-occurrence metric baseline
-Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns (Kandpal et al., 2023; Mallen et al., 2023; Allen-Zhu and Li, 2024b).
-Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach (Kandpal et al., 2023).
-There is a linear relationship between QA accuracy and the logarithm of the number of documents (Kandpal et al., 2023).
+Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns [16, 22, 2].
+Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach [16].
+There is a linear relationship between QA accuracy and the logarithm of the number of documents [16].
 The logarithm of the document count serves as the co-occurrence metric, which constitutes our baseline.
 Knowledge triples
-In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object (Wang et al., 2024).
+In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object [37].
 For a given LLM, denoted as $F$, we define $F$ as mastering the knowledge triple $t$, thus demonstrating competence in the corresponding QA task, as follows:
@@
 Evaluating the knowledge retention of LLMs presents challenges due to the diverse range of queries formed by the subject and relation of knowledge triples.
-Inspired by the evaluation method for fine-tuned LLMs (Ye et al., 2024), we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
+Inspired by the evaluation method for fine-tuned LLMs [40], we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
 Specifically, by constructing synonymous sentences, we generate a substantial number of query templates that are semantically similar but vary in form.
@@
 The co-occurrence metric is given by $\log_{2}(P(s,o))$.
-On the other hand, using the multi-template evaluation method, we obtain the model’s accuracy on QA tasks.
-We fit a linear equation between the co-occurrence metric and QA accuracy, and then calculate R² and Mean Square Error (MSE) to assess the strength and stability of their linear relationship.
+On the other hand, by employing the multi-template evaluation approach, we measure the model’s accuracy on QA tasks.
+To examine the relationship between the co-occurrence metric and QA accuracy, we fit a linear regression model and compute the R² along with the Mean Squared Error (MSE).
+These metrics allow us to evaluate both the strength and the stability of the linear relationship.
 3.2 Considering the Knowledge Specificity to Obtain the MI Metric
@@
 For example, the address of NVIDIA’s headquarters is easier to remember than that of Apple’s headquarters, because the word "Apple" occurs more frequently across diverse contexts.
-MI in information theory considers both the co-occurrence of two variables and their individual specificities (Shannon, 1948).
+MI in information theory considers both the co-occurrence of two variables and their individual specificities [31].
 Therefore, we aim to use MI to improve the co-occurrence metric.
 MI metric
-Unlike the MI in information theory, which considers all possible values of random variables, the MI metric within pre-training data considers only those instances in which the relevant entities actually appear.
-Given pre-training data $\mathbf{P}$ consisting of $\mathbf{N}$ paragraphs, a knowledge triple $(s,r,o)$, and corresponding $P(s)$, $P(o)$, and $P(s,o)$, the MI metric between $s$ and $o$ is defined as:
+[…]
 $I(s,o)$ $=P(s,o)\log\left(\frac{P(s,o)}{P(s)P(o)}\right),$ (2)
@@
 During inference with LLMs, this amount of information refers to the probability of the model generating the object given the subject.
-Considering the uneven distribution of $I(s,o)$, we take its logarithm to obtain the MI metric (Figure 2).
+Considering the uneven distribution of $I(s,o)$ (Figure 2), we take its logarithm to obtain the MI metric.
+Figure 2: Evaluation data statistics. Left: Co-occurrence frequencies of QA entity pairs in The Pile corpus. Right: Distribution of 12,671 QA tasks across 15 relation types.
 3.3 Considering the Model Size to Obtain the Size-dependent MI metric
@@
 The MI metric does not account for the impact of model size.
-Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger (Kaplan et al., 2020).
-In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters (Kandpal et al., 2023).
+Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger [17].
+In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters [16].
 Therefore, we propose to enhance the MI metric by incorporating model size $\Phi$, measured in billions of parameters.
@@
 Since the normalized MI metric is always less than 1, the factor $1+\frac{1}{\Phi}$ decreases as $\Phi$ increases, leading to a larger SMI value, which indicates stronger knowledge retention.
-Figure 2: Evaluation data statistics. Left: Co-occurrence frequencies of QA entity pairs in The Pile corpus. Right: Distribution of 12,671 QA tasks across 15 relation types.
-4 Experiments
-First, we clean and enhance an open-source Wikipedia knowledge triple dataset to make it suitable for inference scenarios using pre-trained models.
-Next, we conduct retrieval analysis on five large-scale open-source pre-trained corpora.
-Finally, we perform experimental analysis on 21 open-source pre-trained models and 3 models that we pre-train ourselves.
-4.1 Experimental Setup
-Evaluation data
-We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours (Elazar et al., 2021).
-We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2.
-For each category of knowledge triple relations, we create 20 synonymous query templates to construct the QA task evaluation dataset.
-Pre-training corpora
-We conduct a document-level retrieval analysis of the following pre-training datasets, selected based on their usage in the LLMs under our consideration:
-  • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs.
-It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality (Gao et al., 2021).
-  • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project.
-It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency (Laurençon et al., 2022; Scao et al., 2022).
-  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs (Soboleva et al., 2023).
-  • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes.
-It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM (Penedo et al., 2023).
-  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency (Foundation, ).
-Pre-training data retrieval
-We process the pre-training corpus at the document level.
-For each knowledge triple, we retrieve the frequency of occurrences of the subject, the object, and their co-occurrences.
-Using the retrieval results from all documents, we calculate the co-occurrence metric and MI metric for each knowledge triple.
-Subsequently, we compute the SMI metric based on the MI metric and the model size.
-The retrieval process runs on 100 CPUs for 3 days.
-Pre-trained models
-We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs (Vaswani et al., 2017).
-These models not only make the model weights available but also provide complete pre-training data.
-They are primarily categorized into the following families:
-  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion (Scao et al., 2022).
-  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile (Black et al., 2021, 2022; Wang, 2021).
-  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters.
-These models use a uniform data and training process, facilitating research that is both replicable and interpretable (Biderman et al., 2023).
-  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata.
-It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions (Zhang et al., 2024; Soboleva et al., 2023; Li et al., 2023).
-The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
-Additionally, we pre-train three models, named ours-1.6b, ours-7b, and ours-13b, based on RefinedWeb and Wikipedia as a supplement to the aforementioned models.
-Evaluation setting
-We assess these pre-trained models using the designated evaluation dataset.
-During the evaluation process, we set the temperature to 0.7, the maximum new tokens to 32, and the precision to bfloat16.
-We consider a response to be correct if the object appears within it.
-To ensure stability in the evaluation, we perform inference 20 times for each of the 20 queries in the QA task, subsequently calculating the average accuracy of 400 responses.
-This inference process runs for 5 days on 8 H20 GPUs.
-The results of the pre-training data retrieval indicate that the co-occurrence frequency distribution of question entities and answer entities is highly uneven (Figure 2).
-The majority of QA data co-occur between 10 and 1000 times.
-In contrast, QA data with higher co-occurrence frequencies are notably sparse.
-To ensure data balance and effectively evaluate data across different ranges, we group QA data with similar metric values.
-For co-occurrence metrics and MI metrics, we calculate the average accuracy using intervals of 0.2.
-The SMI metric is computed based on the MI metric.
-Finally, we perform linear regression on the metrics and accuracy, and calculate the R² to assess the strength of the linear relationship.
-Simultaneously, to evaluate the prediction error of the regression equation in specific QA tasks, we compute the MSE for each QA task and take the average.
+Figure 3: Relationship between average QA accuracy on the full evaluation dataset and model size across five different model families.
+Model size is reported in billions of parameters and plotted on a base-2 logarithmic scale.
+Variations in pre-training data across model families may account for differences in performance.
 Table 1: R² and MSE (×$\times 10^{-2}$) results of the CO-OCCUR, MI, and SMI metrics across different models, ranging from 14M to 176B parameters.
@@
 ours-13b 0.661 0.816 0.820 8.585 7.688 7.674
+4 Experiments
+In the Experimental Setup section, we begin by cleaning and enhancing an open-source Wikipedia knowledge triple dataset to make it suitable for inference with pre-trained models.
+We then introduce five large-scale open-source pre-training corpora and perform retrieval-based analyses on them.
+This is followed by a description of five families of pre-trained language models and the corresponding evaluation protocol.
+In the Experimental Results section, we analyze the performance of 21 open-source pre-trained models along with 3 models that we pre-trained ourselves, comparing them across three evaluation metrics.
+We also conduct a hyperparameter study to further explore model behavior under different configurations.
+4.1 Experimental Setup
+Evaluation data
+We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours [9].
+We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2.
+For each category of knowledge triple relations, we create 20 synonymous query templates to construct the QA task evaluation dataset.
+Pre-training corpora
+We conduct a document-level retrieval analysis of the following pre-training datasets, selected based on their usage in the LLMs under our consideration:
+  • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs.
+It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality [11].
+  • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project.
+It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency [18, 30].
+  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs [32].
+  • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes.
+It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM [26].
+  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency [10].
+Pre-training data retrieval
+We process the pre-training corpus at the document level.
+For each knowledge triple, we retrieve the frequency of occurrences of the subject, the object, and their co-occurrences.
+Using the retrieval results from all documents, we calculate the co-occurrence metric and MI metric for each knowledge triple.
+Subsequently, we compute the SMI metric based on the MI metric and the model size.
+The retrieval process runs on 100 CPUs for 3 days.
+Pre-trained models
+We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs [34].
+These models not only make the model weights available but also provide complete pre-training data.
+They are primarily categorized into the following families:
+  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion [30].
+  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile [4, 5, 35].
+  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters.
+These models use a uniform data and training process, facilitating research that is both replicable and interpretable [3].
+  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata.
+It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions [41, 32, 20].
+The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
+Additionally, we pre-train three models, named ours-1.6b, ours-7b, and ours-13b, based on RefinedWeb and Wikipedia as a supplement to the aforementioned models.
+Evaluation setting
+We assess these pre-trained models using the designated evaluation dataset.
+During the evaluation process, we set the temperature to 0.7, the maximum new tokens to 32, and the precision to bfloat16.
+We consider a response to be correct if the object appears within it.
+To ensure stability in the evaluation, we perform inference 20 times for each of the 20 queries in the QA task, subsequently calculating the average accuracy of 400 responses.
+This inference process runs for 5 days on 8 H20 GPUs.
+The results of the pre-training data retrieval indicate that the co-occurrence frequency distribution of question entities and answer entities is highly uneven (Figure 2).
+The majority of QA data co-occur between 10 and 1000 times.
+To ensure data balance and effectively evaluate data across different ranges, we group QA data with similar metric values.
+For co-occurrence metrics and MI metrics, we calculate the average accuracy using intervals of 0.2.
+The SMI metric is computed based on the MI metric.
+Finally, we perform linear regression on the metrics and accuracy, and calculate the R² to assess the strength of the linear relationship.
+Simultaneously, to evaluate the prediction error of the regression equation in specific QA tasks, we compute the MSE for each QA task and take the average.
+(a) bloom (176b)
+(b) gpt-neox-20b
+(c) pythia-12b
+(d) ours-13b
+Figure 4: Relationship between the SMI metric and average accuracy on QA tasks for different model series: (a) BLOOM (176B), (b) GPT-Neo (20B), (c) Pythia (12B), and (d) Ours (13B).
 (a) $\text{R}^{2}$ on smaller models
@@
 $\uparrow$ Higher is better.
-(a) bloom (176b)
-(b) gpt-neox-20b
-(c) pythia-12b
-(d) ours-13b
-Figure 4: Relationship between the SMI metric and average accuracy on QA tasks for different model series: (a) BLOOM (176B), (b) GPT-Neo (20B), (c) Pythia (12B), and (d) Ours (13B).
 Table 2: Hyperparameter analysis of the SMI metric across different BLOOM models. (a) illustrates the impact of varying unit scales, while (b) shows the effect of different pre-normalization ranges.
@@
 (b) SMI values under different pre-normalization ranges
-Range bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) $[-30,0]$ 0.624 0.756 0.772 0.807 0.814 0.764 $[-35,0]$ 0.626 0.750 0.766 0.800 0.811 0.764 $[-40,0]$ 0.625 0.746 0.761 0.797 0.809 0.764 $[-45,0]$ 0.623 0.742 0.758 0.794 0.808 0.764 $[-50,0]$ 0.620 0.739 0.756 0.793 0.808 0.764
+Range bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) [30,0][-30,0][ - 30 , 0 ] 0.624 0.756 0.772 0.807 0.814 0.764 [35,0][-35,0][ - 35 , 0 ] 0.626 0.750 0.766 0.800 0.811 0.764 [40,0][-40,0][ - 40 , 0 ] 0.625 0.746 0.761 0.797 0.809 0.764 [45,0][-45,0][ - 45 , 0 ] 0.623 0.742 0.758 0.794 0.808 0.764 [50,0][-50,0][ - 50 , 0 ] 0.620 0.739 0.756 0.793 0.808 0.764
 4.2 Experimental Results
 Results of SMI, MI, and co-occurrence metrics
-The primary experimental results for the three metrics are presented in Table 1. Overall, we observe the following: (1) The SMI metric performs
-Figure 5: Relationship between average QA accuracy on the full evaluation dataset and model size across five different model families.
-Model size is reported in billions of parameters and plotted on a base-2 logarithmic scale.
-Variations in pre-training data across model families may account for differences in performance.
-exceptionally well on the evaluation set, with both R² and MSE surpassing the baseline across all models.
+The primary experimental results for the three metrics are presented in Table 1.
+Overall, we observe the following: (1) The SMI metric performs exceptionally well on the evaluation set, with both R² and MSE surpassing the baseline across all models.
 (2) For all models, the SMI metric achieves R² values greater than 0.49, and for models with parameters exceeding 1 billion, the R² exceeds 0.75.
@@
 In contrast, for the BLOOM 176b model, the difference in R² values is only 2.42E-04, and the difference in MSE values is just 6.60E-06.
-This indicates that larger models may possess sufficient memory capacity, thereby diminishing the impact of further scaling.
+This indicates that larger models may possess sufficient memory capacity.
 Moreover, the benefits of data optimization exhibit diminishing returns and an upper limit.
-We find that although the models’ performance on the QA task improves with increasing SMI, the accuracy remains relatively low even when the SMI reaches its maximum value (see Figure 4).
+Although a higher SMI generally correlates with better performance, achieving further gains becomes increasingly difficult at advanced levels due to the logarithmic growth of SMI (Figure 4).
 Specifically, the upper accuracy limits for the models bloom, gpt-neox-20b, pythia-12b, and ours-13b are 0.638, 0.727, 0.75, and 0.771, respectively.
@@
 Although a smaller pre-normalization range is preferable, the minimum value in the experimental data approximates -30, suggesting that further reduction might result in data loss.
+5 Limitations and Future Work
+Although the SMI metric shows strong predictive power for model accuracy on closed-book QA, the current evaluation is limited to this task type.
+This focus is intentional, as closed-book QA serves as a direct probe of factual knowledge internalized during pre-training, rather than general reasoning or transfer ability.
+However, broader task coverage is necessary to assess the general applicability of SMI across diverse task formats.
+Some variation is also observed across model scales, suggesting that future formulations should account for size-invariance.
+In addition, while SMI exhibits empirical effectiveness, its mathematical foundations remain underexplored.
+Future work will aim to establish a more rigorous theoretical grounding by linking information-theoretic principles with LLM behavior.
 6 Conclusion
 In this paper, we explore how to predict model knowledge using only pre-training signals and propose effective solutions to address three major challenges.
-We propose SMI, an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy without any additional training.
-SMI achieves R² > 0.75 on models with over one billion parameters.
-Experimental results further reveal that both scaling model size and optimizing data provide marginal benefits, with the accuracy on specific QA tasks plateauing at approximately 80%.
+We introduce SMI, an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy without any additional training.
+SMI achieves $R^{2}>0.75$ on models with over one billion parameters.
+Experimental results further reveal that both scaling model size and optimizing data yield diminishing returns, with QA accuracy plateauing at around 80% under optimal pre-training conditions.
+This indicates an upper bound on factual knowledge retention during pre-training, highlighting the need for downstream augmentation such as retrieval or prompting.
 This paper emphasizes the feasibility of predicting downstream performance based on pre-training signals.
@@
 To facilitate further research, all evaluation data and analysis code are made publicly available.
-6 Limitations
-Although the SMI metric performs well in predicting the accuracy of pretrained models on specific QA tasks, the prediction results remain highly dependent on the type of task.
-The relationship between the SMI and performance may vary across different categories of tasks, necessitating further investigation into various types of tasks in the future.
-Additionally, the fitting results for models with differing parameter sizes exhibit slight differences, and smaller models demonstrate increased sensitivity to the hyperparameters of the SMI.
-Future research will focus on developing representations of pretrained signals that are more robust to model size and exploring the mathematical foundations underlying these representations.
-Appendix A Technical Appendices and Supplementary Material
-A.1 Licenses and Terms of Use
+Appendix A Licenses and Terms of Use
 We use the following publicly available assets in this work:
 We confirm that all licenses and terms of use are respected.
-A.2 Templates for Knowledge Triples
 Table 3: Twenty templates and their corresponding accuracy values for the $\mathbf{BORN}$ relation on the ours-13b model. [S] represents the subject in the knowledge triple.
@@
 [S]’s starting point was 0.282
-[…]
-A.3 Case Study
 Table 4: Several cases with significant deviations from the co-occurrence metric are shown here.
@@
 Ohio The capital of Ohio is Columbus 11,894,303 3,718,483 1,210,957 0.193
+Appendix B Templates for Knowledge Triples
+[…]
+Appendix C Case Study
 Table 4 highlights four counterintuitive cases in our pre-training data.
@@
 This insight suggests that in downstream task training, even when domain-specific data is significantly fewer than the pre-training data, minimizing repetition with the pre-training data can increase the SMI, making it easier for LLMs to retain the relevant knowledge.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-[…]
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-[…]
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-[…]
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-[…]
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-[…]
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-[…]
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-[…]
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-[…]
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-[…]
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A] This paper constitutes fundamental research and is not associated with any specific applications or deployments.
-[…]
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-[…]
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-[…]
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-[…]
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-[…]
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-[…]
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [Yes]
-[…]
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v5

@@
 Abstract
-The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent.
-Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets.
-In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools.
-We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance.
-In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models.
-We then develop a multi-template QA evaluation framework incorporating paraphrased question variants.
-Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training.
-Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters.
-Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
-Abstract
-The GPT-4 technical report highlights the possibility of predicting model performance on downstream tasks using only pre-training signals, though detailed methodologies are absent.
-Such predictive capabilities are essential for resource-efficient pre-training and the construction of task-aligned datasets.
-In this paper, we aim to predict performance in closed-book question answering (QA), a vital downstream task that directly reflects a model’s internalized knowledge without the help of external tools.
-We address three primary challenges: (1) limited access to and understanding of pre-training corpora, (2) limitations of current evaluation methods for pre-trained models, and (3) limitations of frequency-based metrics in predicting model performance.
-In response, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained large language models.
-We then develop a multi-template QA evaluation framework incorporating paraphrased question variants.
-Building on these foundations, we propose Size-dependent Mutual Information (SMI), an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy, without requiring additional training.
-Experimental results show that SMI outperforms co-occurrence-based baselines, achieving $R^{2}>0.75$ on models with over one billion parameters.
-Theoretical analysis further suggests an upper bound of around 80% QA accuracy under optimal pre-training, reflecting intrinsic memory limitations and motivating the use of retrieval or few-shot methods in later stages.
-[…]
-[…]
-Figure 1: Illustration of the three primary challenges addressed in this work, with the third being the most crucial.
+The GPT-4 technical report suggests that downstream performance can be predicted from pre-training signals, but offers little methodological detail on how to quantify this.
+This work address this gap by modeling knowledge retention, the capacity of a pre-trained language model to memorize factual information from its corpus, and introduce a principled method to estimate it prior to training.
+We propose Size-dependent Mutual Information (SMI), an information-theoretic predictor that integrates knowledge frequency, knowledge specificity, and model size to forecast closed-book question answering (QA) accuracy.
+SMI is validated through large-scale document retrieval over the disclosed pre-training corpora of 21 public and 3 custom models, combined with a robust multi-template QA evaluation.
+Experiments show that SMI significantly outperforms repetition-based baselines and achieves R² > 0.7 in predicting QA accuracy for models above 1B parameters, without additional training.
+The analysis further reveals diminishing returns from scaling data and model size and provides evidence for an intrinsic upper bound on knowledge retention achievable by pre-training alone, motivating retrieval and other augmentation strategies.
+The dataset and code are available at https://github.com/yuhui1038/SMI.
+Beyond Scaling: Measuring and Predicting the Upper Bound of Knowledge Retention in Language Model Pre-Training
+Changhao Jiangthanks: Equal contribution, Ming Zhang11footnotemark: 1, Yifei Cao11footnotemark: 1, Junjie Ye11footnotemark: 1, Xiaoran Fan11footnotemark: 1, Shihan Dou, Zhiheng Xi, Jiajun Sun, Yi Dong, Yujiong Shen, Jingqi Tong, Baoyu Fan, Tao Gui, Qi Zhangthanks: Corresponding author, Xuanjing Huang Fudan University, Shanghai, China IEIT Systems Co.
+Ltd, Shenzhen, China chjiang25@m.fudan.edu.cn, mingzhang23@fudan.edu.cn, qz@fudan.edu.cn
+Figure 1: Overview of our method.
+The framework consists of three stages: (1) Corpus Analysis and Retrieval, where subject and object (question and answer) entities are extracted from pre-training data and their frequencies and co-occurrences are computed; (2) Information-Theoretic Metric Computation, where knowledge frequency, knowledge specificity, and model size are integrated into the size-dependent mutual information (SMI) metric; and (3) Evaluation via Closed-Book QA, where pre-trained language models are assessed using multi-template evaluation, and SMI is compared with empirical QA accuracy to analyze linear relationships and scaling effects.
 1 Introduction
-The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" [25].
-However, the report provides little detail on the methodology behind this prediction approach.
-This technique is critical for the pre-training of large language models (LLMs).
-On the one hand, the cost of a single pre-training run can reach hundreds of millions of dollars, making re-pre-training a highly challenging endeavor [14, 39, 8].
-The prediction technique, which allows for accurate performance forecasting on certain tasks based only on the pre-training signals, helps mitigate the resource waste.
-On the other hand, it can guide the construction of pre-training data that are more aligned with target tasks, optimizing resource usage [12, 21].
-In this paper, we aim to predict performance in closed-book QA, a vital downstream task indicative of a model’s internal knowledge [29].
-This setting is particularly well aligned with the nature of pre-training data and best reflects a model’s ability to answer fact-based questions.
-Meanwhile, we encounter three major challenges (Figure 1):
-Challenge1: Limited access to and understanding of pre-training corpora.
-Currently, few open-source foundation models fully disclose their pre-training datasets.
-Pre-training from scratch is prohibitively expensive, requiring massive data collection and compute.
-For example, Meta reports that pre-training the 405B-parameter Llama 3 model used 15.6 trillion tokens and 16,000 H100 GPUs, with estimated costs over 100 million dollars [8].
-Challenge2: Limitations of current evaluation methods for pre-trained models.
-Given the characteristics of closed-book QA tasks, we can evaluate a model’s accuracy on these tasks to assess the factual knowledge it has internalized during pretraining [29, 36].
-However, most evaluation methods face challenges, such as being overly sensitive to specific in-context examples and having coarse granularity in test data segmentation [24, 33].
-These issues make it difficult to accurately assess whether the pre-trained model has effectively retained specific knowledge.
-In particular, assessing a model’s knowledge using a single query template is unreliable [40].
-Challenge3: Limitations of frequency-based metrics in predicting model performance.
-The model’s success on the target task relies on its ability to acquire world knowledge during pre-training, a process heavily shaped by the data [1, 37].
-Previous research suggests that models tend to acquire knowledge more effectively when it is frequently encountered within the pre-training data [16, 22, 2].
-However, these frequency-based methods does not take into account the specificity of knowledge, and thus only demonstrates general trends.
-For example, despite similar frequencies of occurrence, the model may struggle to pinpoint the location of Apple’s headquarters due to the pervasive nature of the subject.
-In contrast, identifying NVIDIA’s headquarters might be relatively straightforward.
-Furthermore, frequency-based methods do not consider the impact of model size.
-Our work primarily focuses on effectively addressing these three challenges to predict model knowledge from pre-training signals, thereby estimating the model’s performance on downstream tasks without any additional training.
-Specifically:
-First, we conduct large-scale retrieval and semantic analysis across the pre-training corpora of 21 publicly available and 3 custom-trained LLMs.
-With comprehensive access to the full pre-training data, we are able to systematically trace and examine the knowledge that each model was exposed to during pre-training, enabling a detailed understanding of the relationship between pre-training data and downstream performance.
-Second, inspired by the evaluation methods for fine-tuning LLMs [40], we develop a robust multimodal evaluation approach.
-Recognizing that the same knowledge can correspond to various questioning styles, each style allows for the design of corresponding templates.
-For each category of question-answering tasks, we design 20 semantically equivalent but syntactically diverse evaluation templates, enabling a more accurate and reliable assessment of the model’s knowledge retention.
-Finally, we propose SMI, an information-theoretic metric that linearly associates pre-training data characteristics, model size, and QA accuracy without requiring additional training.
-Experimental results on 24 pre-trained models indicate that the SMI metric effectively predicts a model’s knowledge retention.
-Across various model sizes, ranging from 14 million to 176 billion parameters, SMI significantly outperforms the co-occurrence-based baseline.
-For models with more than 1 billion parameters, the coefficient of determination (R²) between SMI and QA accuracy exceeds 0.75, indicating a strong linear relationship.
-Moreover, the comparative analysis across models shows diminishing returns from both scaling model size and optimizing data.
-As model size increases, its marginal impact on QA performance decreases, suggesting that larger capacity alone cannot guarantee continued gains.
-Likewise, while higher SMI values are generally associated with better accuracy, pushing SMI further becomes increasingly difficult due to its logarithmic growth.
-Notably, even in the largest models within each family, QA accuracy consistently plateaus around 80% when SMI is maximized.
-This suggests a fundamental limitation in how much factual knowledge can be internalized during pre-training, highlighting the need for external augmentation methods such as retrieval or prompting in downstream stages.
-In summary, our contributions are threefold:
+The technical report on GPT-4 states, "We registered predictions for GPT-4’s performance on HumanEval before training completed, using only information available prior to training" (OpenAI, 2023).
+Understanding and predicting the generalization ability of large language models (LLMs) has become a key research question in the study of scaling laws (Kaplan et al., 2020).
+Although prior work has established empirical relationships between loss, data size, and model parameters, these results primarily offer coarse-grained predictions of overall performance and do not address what specific factual knowledge is retained from pre-training corpora.
+Recent studies have shown that LLMs can memorize and regurgitate training data (Carlini et al., 2021; Biderman et al., 2023a), but these findings are largely descriptive and do not provide a systematic way to predict retention for new model–corpus pairs.
+Developing principled approaches for this purpose is crucial both for improving our theoretical understanding of model training dynamics and for guiding practical decisions on data curation and compute allocation (Hoffmann et al., 2022; Alabdulmohsin et al., 2022; Ruan et al., ).
+This work addresses this gap by modeling knowledge retention, the capacity of a pre-trained language model to memorize factual information from its corpus, and introduces a principled method to reliably estimate it prior to training.
+Rather than focusing only on aggregate loss curves or perplexity, we take a more fine-grained and systematic view: we ask, given a particular fact in the pre-training corpus, what is the probability that a model will be able to accurately reproduce it in a closed-book setting after training (Petroni et al., 2019)?
+We term this capacity knowledge retention, and formalize it as a quantitative property of a model–corpus pair.
+This framing allows us to reason about model memory at the level of individual facts and categories, and to make meaningful predictions before committing to the cost of full-scale training.
+We propose Size-dependent Mutual Information (SMI), an information-theoretic predictor that integrates knowledge frequency, knowledge specificity, and model size to forecast closed-book question answering (QA) accuracy (Roberts et al., 2020).
+Since pre-trained language models are optimized only for next-token prediction rather than explicit knowledge retrieval, closed-book QA provides a natural and challenging probe for assessing factual recall.
+Our formulation builds on the intuition that both the frequency of a fact in the corpus and its rarity influence its likelihood of being learned, and that these effects are jointly influenced by model size.
+Concretely, SMI captures the mutual information between knowledge subjects and objects (or questions and answers), combining these statistics with model size.
+The resulting predictor is simple to compute from pre-training data and scales gracefully with corpus size, making it suitable for practical use in large-scale settings.
+SMI is validated through large-scale document retrieval over the disclosed pre-training corpora of 21 public and 3 custom models, combined with a robust multi-template QA evaluation.
+Our experimental setup reconstructs the exposure of facts in the training data, then evaluates whether models can correctly answer corresponding questions in a closed-book setting.
+We design multi-template QA prompts to reduce prompt variance and report results across a diverse set of model families, covering a range from 14M to 176B parameters.
+Experiments show that SMI significantly outperforms repetition-based baselines and achieves the coefficient of determination (R²) > 0.7 in predicting QA accuracy for models above 1B parameters, without additional training.
+Relative to heuristics based solely on repetition counts, SMI yields a stronger correlation with observed QA accuracy, particularly in the large-model regime where memorization behavior appears more stable.
+Moreover, we observe that SMI maintains comparable predictive performance across multiple model families and domains, indicating that it captures a generalizable aspect of factual knowledge acquisition in pre-trained language models.
+The analysis further reveals diminishing returns from scaling both data and model size, suggesting that pre-training alone faces an intrinsic limit in its ability to retain factual knowledge.
+Recent studies have similarly observed a performance ceiling on knowledge-intensive tasks, even with substantial scaling of parameters and data (Zhang et al., 2025).
+These observations point to inherent limitations of purely parametric approaches: beyond a certain scale, further increases in model size or training data yield only marginal gains in closed-book performance.
+This motivates exploring hybrid approaches that augment parametric memory with non-parametric mechanisms, including retrieval-augmented generation, external memory modules, or targeted fine-tuning, to mitigate the ceiling effects observed in pre-training.
+Overall, our contributions are threefold:
   1. 1.
-    We propose an information theory-based method and introduce the SMI metric, which accurately predicts the knowledge retention of pre-trained LLMs using only pre-training signals, without requiring any additional training.
+    We formalize knowledge retention as a quantitative metric defined over model–corpus pairs, providing a new way to analyze what models actually memorize during pre-training.
   2. 2.
-    Across LLMs with varying sizes from 14 million to 176 billion parameters, the SMI metric exhibits a strong linear relationship with QA accuracy, significantly outperforms the traditional co-occurrence-based baseline.
+    We propose SMI, a data-driven and training-free predictor that integrates knowledge frequency, knowledge specificity, and model size to estimate closed-book QA accuracy.
   3. 3.
-    Experimental results across diverse models reveal that both scaling model size and optimizing data yield diminishing returns.
-Even under optimal conditions, QA accuracy plateaus around 80%, reflecting the intrinsic limit of knowledge that can be memorized during pre-training.
+    We empirically evaluate SMI on 24 models, showing that it achieves strong predictive accuracy, characterizing diminishing returns from scaling data and model size, and suggesting the presence of an upper bound on knowledge retention achievable by pre-training alone.
 2 Related Work
@@
 For example, Carlini et al.
-[6] identified a logarithmic relationship between repetition and memory effects, while Chowdhery et al.
-[7] showed that sequences repeated over 500 times are completed with over 40% accuracy.
+(2023) identified a logarithmic relationship between repetition and memory effects, while Chowdhery et al.
+(2023) showed that sequences repeated over 500 times are completed with over 40% accuracy.
 Ju et al.
-[15] found that high-frequency data can lead to “factual shortcuts” in multi-hop reasoning.
-Allen-Zhu and Li [2] suggested that exposing knowledge 1000 times enables a storage capacity of two bits per parameter.
+(2024) found that high-frequency data can lead to “factual shortcuts” in multi-hop reasoning.
+Allen-Zhu and Li (2024b) suggested that exposing knowledge 1000 times enables a storage capacity of two bits per parameter.
 Additionally, Razeghi et al.
-[28] and Yadlowsky et al.
-[38] linked lower-order co-occurrences to improved numerical reasoning, while McCoy et al.
-[23] associated task prevalence in pre-training with better performance on tasks like ROT13.
+(2022) and Yadlowsky et al.
+(2023) linked lower-order co-occurrences to improved numerical reasoning, while McCoy et al.
+(2023) associated task prevalence in pre-training with better performance on tasks like ROT13.
 Model size and knowledge retention
 Furthermore, the model size significantly influences the knowledge retention of LLMs.
-Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern [17].
+Research by OpenAI demonstrates that as the number of non-embedding parameters in a model increases, the test loss decreases following a power law pattern (Kaplan et al., 2020).
 Kandpal et al.
-[16] evaluate the BLOOM model families [30] on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters.
-Additionally, the memory capacity of a model closely correlates with its size [19].
+(2023) evaluate the BLOOM model families (Scao et al., 2022) on long-tail QA tasks and find that the fact-based QA accuracy is highly linearly correlated with the logarithm of the number of model parameters.
+Additionally, the memory capacity of a model closely correlates with its size (Li et al., 2024).
 Knowledge triples as an evaluation tool
@@
 Petroni et al.
-[27] employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks.
+(2019) employ the LAMA method to evaluate the implicit knowledge of models like BERT, demonstrating the value of knowledge triples in reasoning tasks.
 Subsequently, He et al.
-[13], Ju et al.
-[15] further explore the application of knowledge triples in reasoning tasks.
+(2024); Ju et al.
+(2024) further explore the application of knowledge triples in reasoning tasks.
 Moreover, Wang et al.
-[37] and Allen-Zhu and Li [1] use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
+(2024) and Allen-Zhu and Li (2024a) use knowledge triples to explore LLMs’ ability to retain and extract knowledge.
 3 Methodology
-To address the three challenges outlined in Section 1, we propose a method to predict model knowledge solely from pre-training signals.
-We begin with prior work and explore the use of the co-occurrence metric.
-Inspired by information theory, we propose the Mutual Information (MI) metric that simultaneously considers both the frequency and specificity of knowledge, significantly enhancing prediction outcomes.
-Finally, we propose the SMI metric, a composite representation of pre-training data characteristics and model size.
+To predict a model’s ability to recall factual knowledge before training, we develop a framework that estimates closed-book QA accuracy directly from pre-training data statistics and model size.
+We first revisit prior work and evaluate a simple repetition-based co-occurrence metric as a baseline.
+Building on information-theoretic principles, we then introduce the Mutual Information (MI) metric, which accounts for both the frequency and specificity of knowledge items and yields improved predictive performance.
+Finally, we extend MI to SMI, which incorporates model size as a scaling factor, providing a more accurate and scalable predictor of knowledge retention.
 3.1 Starting from the Co-occurrence Metric
 Co-occurrence metric baseline
-Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns [16, 22, 2].
-Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach [16].
-There is a linear relationship between QA accuracy and the logarithm of the number of documents [16].
+Prior work indicates that the more frequently knowledge appears in pre-training data, the better the model learns (Kandpal et al., 2023; Mallen et al., 2023; Allen-Zhu and Li, 2024b).
+Specifically, linking the entities in the questions and answers of a QA task to the pre-training data and counting the number of documents that jointly link the question and answer entities is a feasible approach (Kandpal et al., 2023).
+There is a linear relationship between QA accuracy and the logarithm of the number of documents (Carlini et al., 2023; Kandpal et al., 2023).
 The logarithm of the document count serves as the co-occurrence metric, which constitutes our baseline.
 Knowledge triples
-In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object [37].
+In QA tasks, entity relationships can be represented as knowledge triples, denoted by tuples $t=(s,r,o)$, where $s$ is the subject, $r$ is the relation, and $o$ is the object (Wang et al., 2024).
 For a given LLM, denoted as $F$, we define $F$ as mastering the knowledge triple $t$, thus demonstrating competence in the corresponding QA task, as follows:
@@
 Evaluating the knowledge retention of LLMs presents challenges due to the diverse range of queries formed by the subject and relation of knowledge triples.
-Inspired by the evaluation method for fine-tuned LLMs [40], we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
+Inspired by the evaluation method for fine-tuned LLMs (Ye et al., 2024), we develop a robust multi-template evaluation method to assess the accuracy of pre-trained models in QA tasks.
 Specifically, by constructing synonymous sentences, we generate a substantial number of query templates that are semantically similar but vary in form.
@@
 We perform document-level retrieval on the pre-training data and compute three values: $P(s)$ represents the proportion of documents containing the subject out of the total number of documents $\mathbf{N}$ in the corpus, $P(o)$ denotes the proportion of documents containing the object out of $\mathbf{N}$, and $P(s,o)$ signifies the proportion of documents containing both the subject and the object out of $\mathbf{N}$.
-The co-occurrence metric is given by $\log_{2}(P(s,o))$.
+The co-occurrence metric is defined as $\log(P(s,o))$, representing the logarithm of the fraction of documents in which both the subject and object appear.
+For consistency with information-theoretic conventions (Shannon, 1948a), all logarithms in this paper use base 2.
+Since specifying the subject and object typically determines the relation, we omit the relation component for simplicity.
 On the other hand, by employing the multi-template evaluation approach, we measure the model’s accuracy on QA tasks.
@@
 For example, the address of NVIDIA’s headquarters is easier to remember than that of Apple’s headquarters, because the word "Apple" occurs more frequently across diverse contexts.
-MI in information theory considers both the co-occurrence of two variables and their individual specificities [31].
+MI in information theory considers both the co-occurrence of two variables and their individual specificities (Shannon, 1948b).
 Therefore, we aim to use MI to improve the co-occurrence metric.
 MI metric
-[…]
+Unlike the MI in information theory, which considers all possible values of random variables, the MI metric within pre-training data considers only those instances in which the relevant entities actually appear.
+Given pre-training data $\mathbf{P}$ consisting of $\mathbf{N}$ paragraphs, a knowledge triple $(s,r,o)$, and corresponding $P(s)$, $P(o)$, and $P(s,o)$, the MI metric between $s$ and $o$ is defined as:
 $I(s,o)$ $=P(s,o)\log\left(\frac{P(s,o)}{P(s)P(o)}\right),$ (2)
-$MI(s,o)$ $=\log_{2}(I(s,o)).$ (3)
+$MI(s,o)$ $=Norm(\log(I(s,o))).$ (3)
 In the above formula, $P(s)$ and $P(o)$ penalize the number of documents containing the subject and object, respectively, to emphasize the specificity of $s$ and $o$.
@@
 During inference with LLMs, this amount of information refers to the probability of the model generating the object given the subject.
-Considering the uneven distribution of $I(s,o)$ (Figure 2), we take its logarithm to obtain the MI metric.
+Considering the uneven distribution of $I(s,o)$ (see Figure 2), we apply a logarithmic transformation followed by normalization to map its values to the range $[0,1]$, resulting in the MI metric.
+In rare cases where $I(s,o)$ becomes zero or negative due to data sparsity or noise, such instances are regarded as anomalies and excluded from further analysis.
 Figure 2: Evaluation data statistics. Left: Co-occurrence frequencies of QA entity pairs in The Pile corpus. Right: Distribution of 12,671 QA tasks across 15 relation types.
@@
 The MI metric does not account for the impact of model size.
-Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger [17].
-In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters [16].
-Therefore, we propose to enhance the MI metric by incorporating model size $\Phi$, measured in billions of parameters.
+Existing research indicates that the capabilities of LLMs increase with model size, but the returns diminish as the size becomes larger (Kaplan et al., 2020).
+In long-tail QA tasks, the QA accuracy is highly linearly correlated with the logarithm of the number of model parameters (Kandpal et al., 2023).
+Therefore, we propose to extend the MI metric by explicitly incorporating the parameter count $\Phi$ as a measure of model size.
 The method of integrating model size is not unique.
@@
 Size-dependent MI metric
-Based on these principles, we propose the SMI metric—an exponential function with the normalized MI as the base and $1+\frac{1}{\Phi}$ as the exponent.
-$SMI(s,o,\Phi)=Norm(MI(s,o))^{1+\frac{1}{\Phi}}.$ (4)
+Since linear transformations do not alter the strength of linear relationships, an exponential formulation offers a simple yet effective way to incorporate model size while capturing non-linear effects.
+Based on these principles, we propose the SMI metric, an exponential function with the MI metric as the base and $1+\frac{1}{\log\Phi}$ as the exponent.
+$SMI(s,o,\Phi)=MI(s,o)^{1+\frac{1}{\log\Phi}}.$ (4)
 In the above formula, the MI metric quantifies the information of knowledge triples in the pre-training data, while the model size $\Phi$ implies the memory capacity of LLMs.
 These two factors jointly determine the model’s knowledge retention.
-We normalize the MI metric from its original range of -30 to 0 onto a 0–1 scale.
-Since the normalized MI metric is always less than 1, the factor $1+\frac{1}{\Phi}$ decreases as $\Phi$ increases, leading to a larger SMI value, which indicates stronger knowledge retention.
-Figure 3: Relationship between average QA accuracy on the full evaluation dataset and model size across five different model families.
-Model size is reported in billions of parameters and plotted on a base-2 logarithmic scale.
-Variations in pre-training data across model families may account for differences in performance.
-Table 1: R² and MSE (×$\times 10^{-2}$) results of the CO-OCCUR, MI, and SMI metrics across different models, ranging from 14M to 176B parameters.
-Across all models, MI consistently outperforms CO-OCCUR, and SMI further improves upon MI.
-Model R² ($\uparrow$ higher is better) MSE (×$\times 10^{-2}$, $\downarrow$ lower is better)
-CO-OCCUR MI SMI CO-OCCUR MI SMI
-BLOOM
-bloom-560m 0.384 0.589 0.624 8.566 7.729 7.652
-bloom-1b1 0.577 0.715 0.756 9.123 7.950 7.824
-bloom-1b7 0.584 0.740 0.772 9.389 8.083 7.970
-bloom-3b 0.705 0.782 0.807 9.314 7.863 7.764
-bloom-7b1 0.676 0.804 0.814 9.943 8.309 8.263
-bloom (176b) 0.644 0.764 0.764 7.327 6.029 6.029
-GPT-Neo
-gpt-neo-125m 0.349 0.531 0.593 7.004 7.107 6.999
-gpt-neo-1.3B 0.671 0.810 0.859 8.233 7.444 7.265
-gpt-neo-2.7B 0.672 0.826 0.850 8.168 7.320 7.233
-gpt-j-6b 0.631 0.862 0.869 9.026 7.720 7.692
-gpt-neox-20b 0.593 0.841 0.843 8.557 7.476 7.468
-Pythia
-pythia-14m 0.240 0.375 0.496 6.855 7.183 6.978
-pythia-31m 0.287 0.381 0.549 6.905 7.202 6.905
-pythia-70m 0.355 0.430 0.585 6.749 7.051 6.783
-pythia-160m 0.454 0.522 0.611 6.877 6.969 6.796
-pythia-410m 0.524 0.701 0.759 7.439 7.035 6.889
-pythia-1.4b 0.574 0.799 0.835 8.254 7.325 7.202
-pythia-2.8b 0.631 0.845 0.866 8.593 7.505 7.421
-pythia-6.9b 0.637 0.855 0.865 8.538 7.373 7.334
-pythia-12b 0.602 0.853 0.858 8.741 7.514 7.494
-TinyLlama
-TinyLlama-1.1B 0.641 0.760 0.839 7.803 7.135 6.887
-Ours
-ours-1.6b 0.710 0.794 0.848 8.563 7.696 7.462
-ours-7b 0.784 0.847 0.855 8.983 8.033 7.997
-ours-13b 0.661 0.816 0.820 8.585 7.688 7.674
+Since the MI metric is always less than 1, the factor $1+\frac{1}{\Phi}$ decreases as $\Phi$ increases, leading to a larger SMI value, which indicates stronger knowledge retention.
 4 Experiments
@@
 This is followed by a description of five families of pre-trained language models and the corresponding evaluation protocol.
-In the Experimental Results section, we analyze the performance of 21 open-source pre-trained models along with 3 models that we pre-trained ourselves, comparing them across three evaluation metrics.
-We also conduct a hyperparameter study to further explore model behavior under different configurations.
+In the Experimental Results section, we analyze the performance of 21 open-source pre-trained models along with 3 models that we pre-trained ourselves, comparing them across metrics.
 4.1 Experimental Setup
 Evaluation data
-We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours [9].
+We make use of the high-quality knowledge triple dataset ParaRel, which is designed to evaluate whether language models align with factual knowledge, a task similar to ours (Elazar et al., 2021).
 We select 15 categories of knowledge triple relations with a larger quantity of data and clearer semantics, totaling 12,671 knowledge triples, as shown in Figure 2.
@@
   • The Pile: This is an open-source, large-scale text dataset specifically designed for training LLMs.
-It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality [11].
+It encompasses a variety of sources, including webpages, academic papers, code, and books, with an emphasis on diversity and high quality (Gao et al., 2021).
   • ROOTS (En): This dataset is the English subset of ROOTS, a comprehensive multilingual corpus constructed by the BIGScience project.
-It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency [18, 30].
-  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs [32].
+It is used for training the BLOOM models and emphasizes linguistic diversity and data transparency (Laurençon et al., 2022; Scao et al., 2022).
+  • SlimPajama: Constructed by Cerebras, this open-source dataset is a large-scale text corpus aimed at providing high-quality training materials for LLMs (Soboleva et al., 2023).
   • RefinedWeb: Developed by the Abu Dhabi Technology Innovation Institute (TII), this large-scale English web dataset is based on CommonCrawl and undergoes rigorous content extraction, filtering, and deduplication processes.
-It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM [26].
-  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency [10].
+It is intended to serve as high-quality pre-training material for LLMs such as Falcon LLM (Penedo et al., 2023).
+  • Wikipedia: This structured text corpus contains Wikipedia articles and is extensively used for language model training and evaluation due to its high-quality content, comprehensive coverage, and semantic consistency (Foundation, ).
 Pre-training data retrieval
@@
 Subsequently, we compute the SMI metric based on the MI metric and the model size.
-The retrieval process runs on 100 CPUs for 3 days.
-Pre-trained models
-We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs [34].
-These models not only make the model weights available but also provide complete pre-training data.
-They are primarily categorized into the following families:
-  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion [30].
-  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile [4, 5, 35].
-  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters.
-These models use a uniform data and training process, facilitating research that is both replicable and interpretable [3].
-  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata.
-It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions [41, 32, 20].
-The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
-Additionally, we pre-train three models, named ours-1.6b, ours-7b, and ours-13b, based on RefinedWeb and Wikipedia as a supplement to the aforementioned models.
-Evaluation setting
-We assess these pre-trained models using the designated evaluation dataset.
-During the evaluation process, we set the temperature to 0.7, the maximum new tokens to 32, and the precision to bfloat16.
-We consider a response to be correct if the object appears within it.
-To ensure stability in the evaluation, we perform inference 20 times for each of the 20 queries in the QA task, subsequently calculating the average accuracy of 400 responses.
-This inference process runs for 5 days on 8 H20 GPUs.
-The results of the pre-training data retrieval indicate that the co-occurrence frequency distribution of question entities and answer entities is highly uneven (Figure 2).
-The majority of QA data co-occur between 10 and 1000 times.
-To ensure data balance and effectively evaluate data across different ranges, we group QA data with similar metric values.
-For co-occurrence metrics and MI metrics, we calculate the average accuracy using intervals of 0.2.
-The SMI metric is computed based on the MI metric.
-Finally, we perform linear regression on the metrics and accuracy, and calculate the R² to assess the strength of the linear relationship.
-Simultaneously, to evaluate the prediction error of the regression equation in specific QA tasks, we compute the MSE for each QA task and take the average.
+The retrieval process runs on 20 CPUs for 15 days.
+Table 1: R² and MSE ($\times 10^{-2}$) results of the CO-OCCUR, MI, and SMI metrics across different models, ranging from 14M to 176B parameters.
+Across all models, MI consistently outperforms CO-OCCUR, and SMI further improves upon MI.
+Model R² ($\uparrow$ higher is better) MSE ($\times 10^{-2}$, $\downarrow$ lower is better)
+CO-OCCUR MI SMI CO-OCCUR MI SMI
+BLOOM
+bloom-560m 0.384 0.589 0.591 8.566 7.729 7.723
+bloom-1b1 0.577 0.715 0.719 9.123 7.950 7.939
+bloom-1b7 0.584 0.740 0.743 9.389 8.083 8.072
+bloom-3b 0.705 0.782 0.786 9.314 7.863 7.849
+bloom-7b1 0.676 0.804 0.807 9.943 8.309 8.296
+bloom 0.644 0.764 0.766 7.327 6.029 6.026
+GPT-Neo
+gpt-neo-125m 0.349 0.531 0.535 7.004 7.107 7.100
+gpt-neo-1.3B 0.671 0.810 0.814 8.233 7.444 7.427
+gpt-neo-2.7B 0.672 0.826 0.829 8.168 7.320 7.307
+gpt-j-6b 0.631 0.862 0.865 9.026 7.720 7.710
+gpt-neox-20b 0.593 0.841 0.843 8.557 7.476 7.469
+Pythia
+pythia-14m 0.240 0.375 0.378 6.855 7.183 7.177
+pythia-31m 0.287 0.381 0.384 6.905 7.202 7.196
+pythia-70m 0.355 0.430 0.434 6.749 7.051 7.044
+pythia-160m 0.454 0.522 0.526 6.877 6.969 6.961
+pythia-410m 0.524 0.701 0.705 7.439 7.035 7.024
+pythia-1.4b 0.574 0.799 0.803 8.254 7.325 7.312
+pythia-2.8b 0.631 0.845 0.848 8.593 7.505 7.491
+pythia-6.9b 0.637 0.855 0.859 8.538 7.373 7.360
+pythia-12b 0.602 0.853 0.856 8.741 7.514 7.503
+TinyLlama
+TinyLlama-1.1B 0.641 0.762 0.767 7.802 7.081 7.065
+Ours
+ours-1.6b 0.710 0.778 0.782 8.563 7.712 7.694
+ours-7b 0.784 0.837 0.839 8.982 8.028 8.016
+ours-13b 0.661 0.807 0.810 8.584 7.674 7.665
 (a) bloom (176b)
@@
 (d) ours-13b
-Figure 4: Relationship between the SMI metric and average accuracy on QA tasks for different model series: (a) BLOOM (176B), (b) GPT-Neo (20B), (c) Pythia (12B), and (d) Ours (13B).
-(a) $\text{R}^{2}$ on smaller models
-(b) $\text{R}^{2}$ on larger models
-Figure 5: Comparison of $\text{R}^{2}$ results for three metrics—CO-OCCUR, MI, and SMI—across models of different sizes.
-(a) Smaller models (bloom-560m, gpt-neo-1.3B, pythia-14m, ours-1.6b); (b) Larger models (bloom, gpt-neox-20b, pythia-12b, ours-13b).
-$\uparrow$ Higher is better.
-Table 2: Hyperparameter analysis of the SMI metric across different BLOOM models. (a) illustrates the impact of varying unit scales, while (b) shows the effect of different pre-normalization ranges.
-(a) SMI values across varying unit scales
-Unit bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) $10^{8}$ 0.590 0.716 0.741 0.783 0.804 0.764 $10^{9}$ 0.600 0.723 0.745 0.785 0.805 0.764 $10^{10}$ 0.624 0.756 0.772 0.807 0.814 0.764 $10^{11}$ 0.394 0.497 0.586 0.760 0.826 0.766 $10^{12}$ 0.343 0.222 0.197 0.235 0.349 0.769
-(b) SMI values under different pre-normalization ranges
-Range bloom-560m bloom-1b1 bloom-1b7 bloom-3b bloom-7b1 bloom (176b) [30,0][-30,0][ - 30 , 0 ] 0.624 0.756 0.772 0.807 0.814 0.764 [35,0][-35,0][ - 35 , 0 ] 0.626 0.750 0.766 0.800 0.811 0.764 [40,0][-40,0][ - 40 , 0 ] 0.625 0.746 0.761 0.797 0.809 0.764 [45,0][-45,0][ - 45 , 0 ] 0.623 0.742 0.758 0.794 0.808 0.764 [50,0][-50,0][ - 50 , 0 ] 0.620 0.739 0.756 0.793 0.808 0.764
+Figure 3: Relationship between the SMI metric and accuracy on QA tasks for different model series: (a) BLOOM (176B), (b) GPT-Neo (20B), (c) Pythia (12B), and (d) Ours (13B).
+Figure 4: Relationship between average QA accuracy on the full evaluation dataset and model size across five different model families.
+Pre-trained models
+We conduct experiments on 21 fully open-source pre-trained Transformer decoder-only pre-trained LLMs (Vaswani et al., 2017).
+These models not only make the model weights available but also provide complete pre-training data.
+They are primarily categorized into the following families:
+  • BLOOM: Created by BigScience, these open-source multilingual LLMs are trained on ROOTS and supports 46 natural languages and 13 programming languages, with parameter sizes ranging from 560 million to 176 billion (Scao et al., 2022).
+  • GPT-Neo: This family includes the GPT-Neo, GPT-J, and GPT-NeoX models, ranging from 125 million to 20 billion parameters, all trained on The Pile (Black et al., 2021, 2022; Wang, 2021).
+  • Pythia: Developed by EleutherAI, this fully open-source series of LLMs ranges from 14 million to 12 billion parameters.
+These models use a uniform data and training process, facilitating research that is both replicable and interpretable (Biderman et al., 2023b).
+  • TinyLlama: Originating from the Singapore University of Technology and Design (SUTD), this open-source LLM is trained on SlimPajama and Starcoderdata.
+It has 1.1 billion parameters and is designed with the Llama 2 architecture and tokenizer to provide efficient and lightweight LLM solutions (Zhang et al., 2024; Soboleva et al., 2023; Li et al., 2023).
+The specific model we utilize is the TinyLlama-1.1B-intermediate-step-480k-1T, which is trained for approximately one epoch.
+Additionally, we pre-train three models, named ours-1.6b, ours-7b, and ours-13b, based on RefinedWeb and Wikipedia as a supplement to the aforementioned models.
+Evaluation setting
+We evaluate these pre-trained models on the designated QA evaluation dataset.
+During inference, we configure the generation parameters with a temperature of 0.7, a maximum of 32 new tokens, and bfloat16 precision.
+A model response is considered correct if the expected object entity appears anywhere within the generated text.
+To obtain stable and reliable estimates of performance, we repeat the inference procedure 20 times for each of the 20 queries in the QA task, resulting in a total of 400 responses per model.
+The final accuracy is computed as the mean over all 400 responses.
+The full evaluation runs for approximately five days on 8 H100 GPUs.
+The retrieval results from the pre-training corpus show that the co-occurrence frequency distribution of question and answer entities is highly skewed (see Figure 2).
+Most QA pairs co-occur between 10 and 1000 times, while only a small portion appears at very low or very high frequencies.
+To mitigate this imbalance and enable meaningful evaluation across different frequency ranges, we group QA data with similar metric values.
+For the co-occurrence and MI metrics, we compute average accuracy by binning values into intervals of 0.2 on the unnormalized scale.
+The SMI metric is then derived from the MI metric after binning.
+Finally, we perform linear regression on the metrics and accuracy, and calculate the R² to assess the strength of the linear relationship.
+Simultaneously, to evaluate the prediction error of the regression equation in specific QA tasks, we compute the MSE for each QA task and take the average.
 4.2 Experimental Results
-Results of SMI, MI, and co-occurrence metrics
-The primary experimental results for the three metrics are presented in Table 1.
-Overall, we observe the following: (1) The SMI metric performs exceptionally well on the evaluation set, with both R² and MSE surpassing the baseline across all models.
-(2) For all models, the SMI metric achieves R² values greater than 0.49, and for models with parameters exceeding 1 billion, the R² exceeds 0.75.
-This highlights a robust linear relationship and effectively predicts QA accuracy.
-(3) The MI metric outperforms the baseline co-occurrence metric, while the SMI metric further exceeds the MI metric, demonstrating the effectiveness of our iterative improvement approach.
-The comparison of experimental results from models of different sizes indicates diminishing returns when increasing model size.
-A linear relationship is observed between the models’ average accuracy on the full evaluation dataset and the logarithm of model size, suggesting that significant gains in QA performance require exponential increases in model size (Figure 3).
-Meanwhile, the advantage of the SMI metric over the MI metric is more pronounced in smaller models (Figure 5).
-As model size increases, this advantage becomes less evident.
-For the BLOOM 560m model, the R² values differ by 3.55E-02, and the MSE values differ by 7.75E-04.
-In contrast, for the BLOOM 176b model, the difference in R² values is only 2.42E-04, and the difference in MSE values is just 6.60E-06.
-This indicates that larger models may possess sufficient memory capacity.
-Moreover, the benefits of data optimization exhibit diminishing returns and an upper limit.
-Although a higher SMI generally correlates with better performance, achieving further gains becomes increasingly difficult at advanced levels due to the logarithmic growth of SMI (Figure 4).
-Specifically, the upper accuracy limits for the models bloom, gpt-neox-20b, pythia-12b, and ours-13b are 0.638, 0.727, 0.75, and 0.771, respectively.
-This indicates that even with perfect optimization of the pre-training data for these models, the accuracy on the QA task cannot exceed an approximate upper limit of 0.8.
-Results of hyperparameter influence on the SMI metric
-The SMI framework involves two critical hyperparameters.
-The first is the unit for model parameter quantity, which is set by default to 1e9.
-The second is the pre-normalization range for the SMI metric, with a default setting of -30 to 0.
-We conduct ablation experiments on these two hyperparameters, and the results are presented in Table 2.
-From the experimental results, we observe the following: (1) The unit for model parameter quantity significantly affects the SMI results.
-Using 1e10 as the unit yields the best overall performance.
-However, as the model size increases, the optimal unit value also increases.
-Smaller models exhibit lower robustness to parameter variations, whereas larger models display enhanced robustness.
-(2) The pre-normalization range has a relatively minor effect on the SMI results.
-The range of -30 to 0 performs best overall.
-Although a smaller pre-normalization range is preferable, the minimum value in the experimental data approximates -30, suggesting that further reduction might result in data loss.
-5 Limitations and Future Work
-Although the SMI metric shows strong predictive power for model accuracy on closed-book QA, the current evaluation is limited to this task type.
-This focus is intentional, as closed-book QA serves as a direct probe of factual knowledge internalized during pre-training, rather than general reasoning or transfer ability.
-However, broader task coverage is necessary to assess the general applicability of SMI across diverse task formats.
-Some variation is also observed across model scales, suggesting that future formulations should account for size-invariance.
-In addition, while SMI exhibits empirical effectiveness, its mathematical foundations remain underexplored.
-Future work will aim to establish a more rigorous theoretical grounding by linking information-theoretic principles with LLM behavior.
+Comparison of SMI, MI, and CO-OCCUR
+Table 1 presents the R² and MSE results of CO-OCCUR, MI, and SMI across 24 models, with sizes ranging from 14M to 176B parameters.
+Overall, MI consistently improves over the co-occurrence baseline.
+For example, on gpt-neo-2.7B, R² increases from 0.672 (CO-OCCUR) to 0.829 (MI), while MSE decreases from 8.168 to 7.307.
+This suggests that jointly modeling frequency and specificity provides more informative signals for predicting closed-book QA performance.
+Incorporating model size into the metric (SMI) yields further gains across all models, though the magnitude of improvement is typically moderate.
+For instance, on pythia-12b, R² rises from 0.853 (MI) to 0.856 (SMI), and on bloom, from 0.764 to 0.766.
+These results suggest that incorporating model size into the metric provides a more principled account of the variance in observed QA accuracy.
+Diminishing returns from data scaling
+Despite the improved predictive power of SMI, we find that knowledge retention is limited by the nature of pre-training data.
+As shown in Figure 3, even when SMI reaches its maximum value of $1$, the closed-book QA accuracy does not exceed moderate levels.
+Specifically, the empirical upper bounds of accuracy for bloom, gpt-neox-20b, pythia-12b, and ours-13b are approximately $0.527$, $0.578$, $0.588$, and $0.630$, respectively.
+These results indicate that simply scaling the data cannot indefinitely improve factual recall.
+Diminishing returns from model scaling
+We further analyze the impact of model size on knowledge retention.
+Figure 4 shows that the average accuracy on the full evaluation set scales approximately linearly with the logarithm of model parameters, which is consistent with prior scaling-law observations.
+However, the growth rate becomes increasingly shallow for very large models: even at 176176B parameters, the accuracy remains below $0.2$.
+This finding suggests that model scaling alone is insufficient for achieving high factual coverage and further motivates the exploration of complementary approaches, such as retrieval-augmented generation or external memory modules.
 5 Conclusion
-In this paper, we explore how to predict model knowledge using only pre-training signals and propose effective solutions to address three major challenges.
-We introduce SMI, an information-theoretic metric that linearly correlates pre-training data characteristics, model size, and QA accuracy without any additional training.
-SMI achieves $R^{2}>0.75$ on models with over one billion parameters.
-Experimental results further reveal that both scaling model size and optimizing data yield diminishing returns, with QA accuracy plateauing at around 80% under optimal pre-training conditions.
-This indicates an upper bound on factual knowledge retention during pre-training, highlighting the need for downstream augmentation such as retrieval or prompting.
-This paper emphasizes the feasibility of predicting downstream performance based on pre-training signals.
-It offers a principled metric for data-model-task alignment, thereby providing a reference for optimizing data selection and performance prediction during the pre-training phase of LLMs.
-To facilitate further research, all evaluation data and analysis code are made publicly available.
-Appendix A Licenses and Terms of Use
+In this work, we introduce a framework for quantifying knowledge retention in pre-trained language models, defined as the probability that a model reproduces factual information from its training corpus in a closed-book setting.
+We propose SMI, an information-theoretic predictor that integrates knowledge frequency, knowledge specificity, and model size, and demonstrate its effectiveness in forecasting closed-book QA accuracy.
+Our large-scale evaluation across 24 models shows that SMI achieves high predictive correlation (R² > 0.7 for models above 1B parameters) and remains robust across model families and domains.
+Beyond prediction, our analysis reveals diminishing returns from scaling data and model size, suggesting an upper bound on factual knowledge that can be stored through pre-training alone.
+At the same time, SMI provides a diagnostic lens for data and model design: increasing the frequency of subject–object co-occurrences can strengthen knowledge memorization, filtering out irrelevant material can improve specificity, and scaling model size can further enhance retention, albeit with diminishing returns.
+These insights suggest that complementary strategies such as retrieval-augmented generation, external memory modules, and targeted fine-tuning may be necessary to further improve factual recall in LLMs.
+We hope that our work provides a foundation for more systematic investigations into the interaction between corpus composition, model size, and factual knowledge retention, and informs future efforts to build models that combine parametric and non-parametric memory in a principled manner.
+6 Limitations
+While our study provides an information-theoretic framework for predicting knowledge retention, several limitations remain.
+First, our analysis is restricted to closed-book QA tasks and does not cover open-ended generation, reasoning-intensive tasks, or multi-hop factual queries, which may involve different memorization dynamics.
+Second, we rely on publicly available corpora and approximate document matching, which may introduce noise when measuring fact frequency and specificity.
+Third, our experiments focus on models with up to tens of billions of parameters, leaving open the question of whether the observed trends hold for frontier-scale models with hundreds of billions of parameters.
+Finally, our approach does not address how training dynamics, such as optimization schedules or curriculum learning, may interact with knowledge retention, which we leave for future work.
+Appendix A AI Assistants in Research or Writing
+In preparing this manuscript, AI assistants were employed solely to assist with refining the clarity, style, and readability of certain text segments.
+They were not involved in designing the study, developing or implementing the methodology, collecting or analyzing data, or generating the primary scientific contributions.
+All substantive research decisions, analyses, and conclusions are fully the responsibility of the authors.
+Appendix B Licenses and Terms of Use
 We use the following publicly available assets in this work:
 We confirm that all licenses and terms of use are respected.
-Table 3: Twenty templates and their corresponding accuracy values for the $\mathbf{BORN}$ relation on the ours-13b model. [S] represents the subject in the knowledge triple.
-Template ACC
-[S]’s birth took place in 0.186
-[S] was brought into existence in 0.187
-[S] was born into the world in 0.190
-[S]’s nativity is 0.197
-The place in which [S] was given birth to is 0.201
-The hometown of [S] is 0.203
-The birthplace of [S] is 0.212
-[S] was delivered in 0.215
-[S]’s cradle was in 0.219
-[S]’s entry into the world happened in 0.225
-[S] first saw the light in 0.228
-[S]’s roots are in 0.230
-[S]’s life began in 0.230
-[S] was born and raised in 0.237
-[S]’s origin is 0.238
-[S] entered life in 0.251
-[S] came into the world in 0.255
-[S] was born in 0.255
-[S]’s birth occurred in 0.263
-[S]’s starting point was 0.282
-Table 4: Several cases with significant deviations from the co-occurrence metric are shown here.
-$N_{s}$ is the number of paragraphs in which the subject appears, $N_{o}$ is the number of paragraphs in which the object appears, and $N_{s,o}$ is the number of paragraphs in which they co-occur.
-The first two cases exhibit rare co-occurrence but demonstrate high accuracy, whereas the latter two cases show frequent co-occurrence but exhibit low accuracy.
-Subject Question Object $N_{s}$ $N_{o}$ $N_{s,o}$ $ACC$
-Rare Co-occurrence but High Accuracy
-Csepel SC The headquarters of Csepel SC is in Budapest 398 748,860 172 0.655
-Liis Lemsalu Liis Lemsalu is a citizen of Estonia 307 967,207 160 0.570
-Frequent Co-occurrence but Low Accuracy
-Paris Paris is located in Europe 20,902,571 59,408,155 3,761,610 0.043
-Ohio The capital of Ohio is Columbus 11,894,303 3,718,483 1,210,957 0.193
-Appendix B Templates for Knowledge Triples
-[…]
 Appendix C Case Study
@@
 This insight suggests that in downstream task training, even when domain-specific data is significantly fewer than the pre-training data, minimizing repetition with the pre-training data can increase the SMI, making it easier for LLMs to retain the relevant knowledge.
+Appendix D Templates for Knowledge Triples
+[…]
+{ "plays": [ "[S] plays", "[S] performs", "[S] engages in playing", "[S] is an musician of", "[S] is involved in playing", "[S] is a performer of", "[S] executes", "[S] is a practitioner of", "[S] is an player of", "[S] is an artist of", "The genre that [S] performs is", "[S] captivates audiences by performing", "[S] performs the musical style of", "[S] is known for his performances in", "[S]’s musical expertise lies in", "[S] is recognized for playing", "[S] is a notable performer of", "[S]’s artistry is showcased in", "[S]’s musical inclinations are towards", "The musical category that [S] excels in is" ], "occupation": [ "The occupation of [S] is", "[S] works as an", "[S] is employed in the profession of an", "[S] pursues a profession as an", "[S] is a professional", "[S] is engaged in the profession of an", "[S]’s vocation is an", "[S] is an professional", "[S]’s role in life is that of an", "[S]’s professional identity is that of an", "The job of [S] is", "[S] is employed as an", "[S] has a career as an", "The profession of [S] is", "[S]’s line of work is", "[S] makes her career as an", "The career of [S] is", "[S] has chosen to be an", "[S] earns a living as an", "[S]’s professional life revolves around being an" ], "work": [ "[S] works in the field of", "[S] specializes in the field of", "[S]’s work focuses on", "[S] is engaged in the area of", "[S]’s professional activity lies in", "[S]’s expertise is in", "[S] operates within the field of", "[S]’s research is concentrated in", "[S] works extensively in", "[S] has dedicated his career to", "[S] is involved in the study of", "[S]’s field of study is", "[S] contributes to the discipline of", "[S] practices in the area of", "[S]’s work domain is", "[S] conducts his work in", "[S]’s field of expertise is", "[S]’s career revolves around", "[S] works professionally in", "[S]’s occupation relates to" ], "is-located": [ "[S] is located in", "The fatherland of [S] is in", "[S] can be found in", "[S] lives in", "[S] is situated within", "[S] is based in", "[S] resides in", "The country in which [S] lives is", "[S]’s location is in", "[S] stays in", "[S]’s address is in", "[S] is present in", "[S] spends most of his time in", "[S] belongs to", "[S]’s position is in", "You can find [S] in", "[S] is positioned in", "[S] exists in", "[S] stays primarily in", "The country in which [S] is located is" ], "locate": [ "[S] is located in", "The place in which [S] is held is", "[S] is held in", "The place where [S] takes place is", "[S] takes place in", "[S] is hosted in", "The venue for [S] is", "[S] happens in", "[S] is organized in", "[S] is staged in", "[S] occurs in", "The place for [S] is", "The location of [S] is", "[S] is based in", "[S] is set in", "[S]’s venue is in", "[S] is arranged in", "[S] is situated in", "People can take part in [S] in", "[S]’s setting is in" ], "found": [ "[S] was founded in", "[S] originated in", "[S] was established in", "[S] came into existence in", "[S] was formed in", "[S] had its beginnings in", "[S] started in", "[S] was created in", "[S] was initiated in", "[S] took shape in", "[S] was launched in", "[S] began in", "[S] emerged in", "The foundation of [S] took place in", "[S] originated from", "[S] was set up in", "[S] was first formed in", "[S] came to be in", "[S] had its origin in", "[S]’s formation occurred in" ], "born": [ "[S] was born in", "[S] came into the world in", "The birthplace of [S] is", "[S] entered life in", "[S] was delivered in", "[S]’s origin is", "[S] first saw the light in", "[S]’s birth occurred in", "[S] was brought into existence in", "[S]’s nativity is", "The place in which [S] was given birth to is", "[S] was born and raised in", "[S]’s roots are in", "The hometown of [S] is", "[S]’s cradle was in", "[S]’s entry into the world happened in", "[S]’s birth took place in", "[S] was born into the world in", "[S]’s life began in", "[S]’s starting point was" ], "die": [ "[S] died in", "[S] passed away in", "[S]’s death occurred in", "[S] lost his life in", "[S]’s demise happened in", "[S] perished in", "[S]’s end came in", "[S]’s life ended in", "[S] met his end in", "[S]’s passing was in", "[S]’s death took place in", "[S]’s final moments were in", "[S]’s last breath was in", "[S]’s life was lost in", "[S]’s expiration occurred in", "[S]’s fatality was in", "[S]’s death scene was in", "The place in which [S] passed away is", "[S]’s death location is", "[S]’s death site is" ], "air": [ "[S] was originally aired on", "[S] first aired on", "[S] premiered on", "[S] was initially broadcast on", "[S] debuted on", "[S] was first shown on", "[S] was originally broadcast on", "[S] was first transmitted on", "[S] was first telecast on", "[S] was first screened on", "[S] was first presented on", "[S] had its first airing on", "[S] was first televised on", "[S] was first released on", "[S] was first featured on", "[S] was first displayed on", "[S] was first exhibited on", "[S] was first put on air on", "[S] was first launched on", "[S] was first introduced on" ], "headquarter": [ "The headquarters of [S] is in", "[S]’s headquarters is in", "[S]’s main office is in", "[S]’s central office is in", "[S]’s head office is in", "The corporate office of [S] is in", "[S]’s principal office is in", "[S]’s administrative center is in", "[S]’s base of operations is in", "[S]’s HQ is in", "The headquarters’s location of [S] is in", "[S]’s headquarters office is in", "[S]’s primary office is in", "[S]’s main administrative office is in", "[S]’s central headquarters is in", "[S]’s main base is in", "The main hub of [S] is in", "[S]’s main building is in", "[S]’s headquarters is located in", "The place in which [S]’s headquarters is situated in" ], "capital": [ "[S] is the capital of", "[S] serves as the capital of", "[S] is recognized as the capital of", "[S] stands as the capital of", "[S] is known as the capital of", "[S] functions as the capital of", "[S] is officially the capital of", "[S] holds the title of capital of", "[S] is designated as the capital of", "[S] is the administrative capital of", "[S] is the political capital of", "[S] is the governmental capital of", "[S] is the principal capital of", "[S] is the main capital of", "[S] is the chief capital of", "[S] is the primary capital of", "[S] is the foremost capital of", "The country whose capital is [S] is", "[S] is known to the people as the capital of", "As the capital city, [S] has an extremely important place in" ], "citizen": [ "[S] is a citizen of", "The country where [S] holds citizenship and enjoys benefits is", "The nation that grants [S] the rights and privileges of citizenship is", "[S] is from the country of", "The place of [S]’s birth and upbringing is", "[S] belongs to", "[S]’s homeland is", "[S]’s home country is", "[S] is a resident of", "[S]’s origin is", "The land where [S]’s roots run deep is", "[S]’s place of origin is", "[S]’s country of origin is", "The nation that [S] calls home is", "[S]’s native country is", "[S]’s nationality is from", "The place where [S] was born and raised is", "[S]’s citizenship gives him the rights and privileges of a citizen in", "[S]’s citizenship makes him a part of the country’s community in", "The country that [S] resides in and holds citizenship is" ], "create": [ "[S] was created in", "The birthplace of [S] is", "[S] originated in", "The invention of [S] took place in", "[S] hails from", "The country that [S] was made in is", "[S] was initially developed in", "The country that introduced [S] is", "[S] was crafted in", "The place of origin for [S] is", "[S] was first produced in", "The homeland of [S] is", "The source of [S] is", "[S] was manufactured in", "The origin of [S]’s creation is", "[S] was designed in", "The place where [S] was conceived is", "[S] was made in", "The birth nation of [S] is", "[S] was developed in" ], "play": [ "[S] plays", "[S] is known for playing the", "[S]’s instrument is the", "[S] is a musician who plays the", "[S] is one who has mastered the art of the", "[S] has a gift for playing the", "Few can rival [S]’s prowess on the", "[S] is a true virtuoso of the", "[S]’s music is infused with the soulful sound of the", "Audiences are captivated by [S]’s", "[S]’s musical talents shine through his", "[S] is a skilled", "One can hear [S]’s passion in every", "With every note, [S] demonstrates his mastery of the", "[S] is a", "[S] plays the", "[S] specializes in the", "[S] performs on the", "Music flows through [S] and his", "As for instruments, [S] prefers the" ], "is-locate": [ "[S] is located in", "[S] lies in the", "[S] lies within the boundaries of", "The place where you can find [S] is", "[S] is situated in", "[S] is positioned in", "[S] exists in", "[S] is placed in", "[S] resides in", "The place where [S] can be fully appreciated is", "[S] can be found in", "[S] is in the area of", "[S] is part of the landscape of", "[S] is in the lands of", "The region that holds [S] is", "The location of [S] is", "[S] is in the sector of", "The place that [S] is a part of is", "[S] is in the section of", "[S] is in the region of" ] }